### v3

# A dataset and model for recognition of audiologically relevant environments for hearing aids: AHEAD-DS and YAMNet+

Abstract
Scene recognition of audiologically relevant environments is important for hearing aids; however, it is challenging, in part because of the limitations of existing datasets. Datasets often lack public accessibility, completeness, or audiologically relevant labels, hindering systematic comparison of machine learning models. Deploying these models on resource-constrained edge devices presents another challenge. Our solution is two-fold: we leverage several open source datasets to create AHEAD-DS, a dataset designed for scene recognition of audiologically relevant environments, and introduce YAMNet+, a sound recognition model. AHEAD-DS aims to provide a standardised, publicly available dataset with consistent labels relevant to hearing aids, facilitating model comparison. YAMNet+ is designed for deployment on edge devices like smartphones connected to hearing devices, such as hearing aids and wireless earphones with hearing aid functionality; serving as a baseline model for sound-based scene recognition. YAMNet+ achieved a mean average precision of 0.83 and accuracy of 0.93 on the testing set of AHEAD-DS across fourteen categories of audiologically relevant environments. We found that applying transfer learning from the pretrained YAMNet model was essential. We demonstrated real-time sound-based scene recognition capabilities on edge devices by deploying YAMNet+ to an Android smartphone. Even with a Google Pixel 3 (a phone with modest specifications, released in 2018), the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio. Our website and code https://github.com/Australian-Future-Hearing-Initiative.
Keywords: scene recognition, sound recognition, audiologically relevant labels, machine hearing, computer audition, AHEAD-DS, YAMNet+
1 Introduction
State of the art hearing devices incorporate scene recognition in order to identify audiological environments and apply appropriate processing (filters, amplification, beam forming, etc) to improve a user’s hearing experience [3].
Several machine learning (ML) models have been proposed for scene recognition [24]. However, there are several limitations: ML models are challenging to port to edge devices [4] and datasets have inconsistent labels (As outlined in Table 1).
It is important that models can run on edge devices to preserve privacy by avoiding the transmission of sensitive data and to minimise the cost of data transmission. However, not all models are immediately deployable to devices, such as hearing aids and wireless earphones with hearing aid functionality (henceforth will be collectively referred to as hearing devices), or to portable devices such as smartphones that operate in tandem with hearing devices. Resource constraints and data format compatibility must be overcome for model deployment on the edge [23].
Models are trained and tested using datasets with different scene labels and are thus not directly comparable. These datasets are either not public or incomplete, and general-purpose sound datasets do not have labels that are necessarily relevant to hearing devices [15].
We were motivated to address these problems for several reasons: There is a gap in publicly available standardised datasets for scene recognition of audiologically relevant environments for hearing devices, and an open-source baseline ML model designed for edge devices is currently unavailable. Our ultimate goal is to improve the hearing experience for users of hearing devices. This paper makes the following contributions:
  • A description of a sound dataset named Another HEaring AiD scenes Data Set (AHEAD-DS). This dataset was created for the purpose of scene recognition of audiologically relevant environments for hearing devices, which allows comparison of models using a standardised dataset with consistent labels. The dataset is released under a permissive licence, in a ready-to-use state, divided into training/validation/testing sets and annotated with the ground truth.
  • A description of a ML model called YAMNet+ for sound based scene recognition. The model can act as a baseline model for further development of sound based scene recognition models. The model is released under a permissive licence, including model definitions, training, testing code, and code to convert it into a format deployable to edge devices.
  • We implemented YAMNet+, trained it using AHEAD-DS, and presented benchmark results including precision, recall and inference speed. We have deployed the model to a smartphone to demonstrate it works on an edge device.
This paper is organised as follows. The second section presents an overview of existing datasets and audiologically relevant environment labels. The third section presents an overview of several sound recognition models. The fourth section provides a description of AHEAD-DS and how the dataset was created. The fifth section describes the YAMNet+ sound recognition model. Section six presents the results of training and testing YAMNet+ using AHEAD-DS. Section seven discusses the overall findings of this paper and describes the possible future direction of the research. Section eight draws some conclusions.
2 Scene Recognition Datasets for Hearing Devices
There are several relevant studies (Table 1) involving scene recognition of audiologically relevant environments for hearing devices.
The availability of a standardised dataset helped drive tremendous advancements in computer vision [2]. It is hoped that a dataset for scene recognition for hearing devices will similarly drive advancements. Hence, we sought such a dataset, but noted the following challenges with existing datasets relevant to hearing devices.
  • Datasets may be proprietary, with a restrictive licence and not uploaded to a publicly accessible repository. This was an issue with four studies [33, 8, 11, 39].
  • Data used in the studies were not recorded. Data was processed on-the-fly and raw sound discarded in three studies [14, 18, 5].
  • In one study [15], the dataset available online does not match the description in the published paper.
  • The dataset is a mix of secondary sources, but the specific subset from these sources are unknown. This was an issue for three studies [15, 33, 31].
  • The dataset is a mix of secondary sources but there is ambiguity in the mixing procedure. For example, there may be pseudo random procedures during processing of sound files, which are not clearly explained in the paper. This was an issue with two studies [15, 31].
  • Each study uses datasets with different labels; therefore, they are not easily comparable. This is the case for all studies in Table 1.
Additionally, several large general-purpose datasets exist for sound recognition. Datasets such as AudioSet [9], UrbanSound8K [28], ESC50 [27], and TAU scenes [35] provide extensive training material for ML models. However, labels are either too specific or too general to be used directly for the purpose here. For example, in AudioSet recordings can be labelled bus or truck. This level of granular sound recognition is unnecessary: it is only necessary to recognise and respond to motor vehicle noises to improve the hearing experience. However, data can still be used through transfer learning or labels can be remapped to more broad labels.
Conversely, TAU scenes provide location-based recordings which are too general, such as at airports and metro stations. The sound of a motor vehicle will be similarly disruptive irrespective of whether they occur at either location. However, labels could be modified and made more specific, but this would likely require listening to and re-annotating the recordings.
UrbanSound8K and ESC50 provide extensive data. Both datasets contain very specific labels, such as the sounds from specific animals. Like AudioSet, this level of granularity is unnecessary for the purposes here. However, the data is still useful for transfer learning.
Given these issues, we created our own data set using HEAR-DS as a starting point.
Table 1: This table summarises relevant scene labels for hearing devices from several studies. Summarised are the data licence with notes on the data collection and processing procedures.
Study Labels Data Licence Notes
[14] Quiet, Speech moderate, Speech high, Noise moderate, Noise high, Speech in noise moderate, Speech in noise high Audio data not saved Sound data processed on-the-fly. The classification results were recorded but sound was not. The classification system was a proprietary part of the hearing aids used in the study.
[15] Cocktail party, Interfering speakers, In traffic, Speech in traffic, In vehicle, Speech in vehicle, Music, Speech in music, Quiet indoors, Speech in qu. Indoors, Reverberant env., Speech in reverb. env. , Wind turbulence, Speech in wind turb. Open source Sound dataset derived from CHiME 5 [1], GTZAN [32], the environment sounds recorded by authors of the study are released publicly. Though the number of samples released is not consistent with the description in the study. The code for mixing speech and environment sounds is missing. Dataset is named Hearing Aid Research Data Set for Acoustic Environment Recognition (HEAR-DS).
[33] Music, Noise, Speech with noise, Speech, Silence Mix of open source and proprietary Sound dataset derived from freesound.org, Demand database [30], CSTR VCTK corpus [37], Million Song Dataset [6], LibriSpeech ASR corpus [26] and YouTube. YouTube source is proprietary.
[8] Bus, Subway, Street, Indoor, Car Proprietary Sound dataset recorded by study authors, not released publicly.
[18] Transportation, Physical, Basics, Social, Listening Audio data not saved Sound features and acceleration features are computed and recorded, but sound was not recorded.
[5] Quiet, Speech, Speech in noise, Noise Audio data not saved Sound data was collected and processed on-the-fly and not saved.
[7] Speech in restaurant, Speech in car, Speech in large room Proprietary The test results and labels come from a separate study [34] provided by a hearing aid manufacturer. The data was not publicly available.
[11] Music, Single talker food court, One talker on subway platform, Conversation in quiet Proprietary Labels from five hearing aid manufacturers are mapped to the labels used in the study. Sound was recorded by the authors and not released publicly.
[31] Vehicle engine, Machine tools, Household appliance, Nature, Human speech Open source Sound dataset derived from UrbanSound8K [28] and ESC50 [27]. Remapped to relevant labels. An unknown amount of data from the two source datasets was removed. Data augmentation is used but it was not described in sufficient detail to be reproducible.
[39] Speech in quiet, Speech in noise, Noise, Music Mix of open source and proprietary Sound dataset derived from proprietary Phonak Sound CD 2, self recordings on the London tube subway, and a song from royalty free freesound.org.
3 Sound Recognition Machine Learning Models
Three machine learning models relevant for sound recognition are presented in this section, PANN [17], AST [10], and YAMNet [38]. The models in all of these studies were trained, validated, and tested using the AudioSet dataset. The findings are summarised in Table 2. The extensive training dataset makes any of the models suitable for tuning via transfer learning.
While more sound recognition models exist, they are minor variations on the three aforementioned models [40]. Most models make use of frequency domain based features, such as log mel spectrograms. They use either convolution or attention neural network blocks with dense layers. The spectrograms are treated as images and processed using an image processing neural network. Minor variations involve switching spectrograms with log mel spectrograms, or a different configuration of neural network blocks. A minority of models such as PANN use the raw waveform as a feature.
The three models PANN, AST and YAMNet have the following tradeoffs.
  • PANN and AST achieve better mean average precision (mAP) results compared to YAMNet, 0.44 and 0.46 vs 0.31 respectively, using the testing set of AudioSet. However, PANN and AST require over 20 times more parameters compared to YAMNet, approximately 81 and 86 million compared to 3.7 million respectively. This poses a challenge for deploying PANN and AST to resource-constrained edge devices.
  • YAMNet is implemented using the TensorFlow ML framework, while PANN and AST are implemented using the Pytorch ML framework. A robust software pipeline exists to port any TensorFlow model to an edge device, such as an Android phone, using TensorFlow Lite. However, YAMNet’s implementation is not fully open source, while model definitions and weights have been released, training code is missing from the code repositories.
To address these issues, we developed our own model by creating a variation of YAMNet.
Table 2: This table details several machine learning models for sound recognition. The mAP for the AudioSet testing set and number of parameters are those reported by the authors of each study. PANN and AST have higher mAP and use significantly more parameters compared to YAMNet/YAMNet+. This suggests that while PANN and AST have better sound recognition capabilities, they demand significantly more computational resources.
Model Features Architecture Parameters mAP
YAMNet [38] / YAMNet+ Log mel spectrogram Convolution 3,751,369 0.31
PANN [17] Log mel spectrogram, waveform Convolution 81,065,487 0.44
[…]
4 Another HEaring AiD scenes Data Set (AHEAD-DS)
We leverage several open source datasets to create a dataset which we called AHEAD-DS. A summary of the dataset is shown in Table 3. A summary of the sources of the dataset is shown in Table 4. There are a total of 9968 recordings, 10s per recording, across 14 different labels. Each instance is single-channel, 16-bit integers and sampled at 16 kHz.
Despite the limitations outlined earlier, AHEAD-DS is based on HEAR-DS and CHiME 6 Dev [36]. HEAR-DS provides a solid foundation for the following reasons:
  • The recordings of environment sounds have been open sourced with a permissive licence.
  • According to the authors of HEAR-DS, the 14 labels have been chosen from audiologically relevant environments for hearing devices by audiologists and Audifon, a manufacturer of hearing aids.
AHEAD-DS follows the design of HEAR-DS, it employs an identical set of 14 labels. This set comprises six environmental sound categories and an additional six categories representing speech instances within these respective environments. Hearing devices likely require different signal processing strategies to optimise speech intelligibility in each acoustic setting [3]. Therefore, the detection of speech embedded within specific environments represents a step towards creating an improved hearing experience for a hearing devices user.
The main issue with HEAR-DS is that the dataset is not in a fixed, ready-to-use state. This is because the published data is unmixed, and some pseudo-random procedures were used during the production of the HEAR-DS dataset. Since the code has not been published, these procedures cannot be fully reproduced. This causes the following problems.
  • It is not possible to determine the split for training and testing.
  • It is not possible to produce results which can be compared directly against the ones in the HEAR-DS paper.
The advantage of AHEAD-DS is that the dataset is in a ready-to-use state. The code for the mixing procedure is published and all mixing parameters are documented and reproducible. This is important because it removes a variable when comparing ML models. The speech and environment sounds have been released in both mixed and unmixed forms, eliminating any ambiguity regarding the subsets used to derive our new dataset.
The sounds from each environment for AHEAD-DS were taken from HEAR-DS. The authors of HEAR-DS recorded each scene using a dummy head. The head was equipped with a synthetic ear, the GRAS KB1065/1066 Pinnae. An Audiofon microphone was placed in each ear of the dummy head in an in-the-canal position. A Focusrite Scarlett 18i6 sound interface was used for analogue to digital conversion. All recordings were initially captured as 32-bit floating-point values sampled at 48 kHz, and later downsampled to a 16 kHz sampling rate with 16-bit integer quantisation. There were a total of 5929 clips of 10s, from 6 different environments.
Music samples from 10 genres were taken from HEAR-DS. The authors of HEAR-DS obtained the music from GTZAN [32]. All music samples were 10s long, 16-bit integers and sampled at 16 kHz. There were a total of 2992 clips of 10s of music.
One issue raised about AHEAD-DS and HEAR-DS was the constant acoustic signature of the reverberant environment might interfere with the classification of the acoustic signature of the other environments. For sound recognition, if a human can identify the presence of a reverberant environment, an acoustic signature exists for reverberant environments. Deep learning will automatically generate compensating features; these features are issues of concern for manual feature engineering; but not for deep learning [19].
CHiME 6 Dev was used for speech samples in AHEAD-DS instead of CHiME 5, which was used in HEAR-DS. Both contain the same speech recordings, but the start and end times of all recordings have been corrected in CHiME 6 Dev to ensure all recordings are synchronised. CHiME 6 dev is a subset of CHiME 6; it is smaller than the whole dataset but sufficiently long for mixing with environment sounds to provide all the speech in $\langle environment\rangle$ clips and to produce interfering speakers. The interfering speakers clips were derived entirely from CHiME 6 dev.
The authors of CHiME 6 dev recorded the speech in two sessions. Each session in two different home environments consisting of an open space with dining, kitchen, and lounge areas. A total 44 sound files from CHiME 6 Dev were used in AHEAD-DS, 24 recordings from the first environment and 20 in the second. Each session consisted of four people and would begin by having participants take turns to introduce themselves via speaking and then read out a short passage. Once this is complete, participants engage in casual conversation as they prepare and share a meal together.
The authors of CHiME 6 dev made the recordings by placing six Microsoft Kinects throughout the home in the first session and five in the second. However, the Kinect version is unspecified. Each Kinect recorded four sound streams simultaneously, using 16-bit integers, sampled at 16 kHz.
In CHiME 6 additional speech was recorded using body-worn microphones, but only the recordings from the Kinect were used. The recordings from body-worn microphones lacked acoustic realism as they sounded like clean studio recordings.
The label interfering speakers contains only speech taken from CHiME 6 and is not mixed with environment sounds. It represents a scenario with multiple speakers where the speech is intelligible. The label cocktail party contains multiple speakers but the speech is unintelligible.
Compared to datasets like AudioSet, AHEAD-DS is still relatively small, which is a disadvantage. In later sections, we analyse the implications of having a small dataset (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
Table 3: This table summarises the contents of AHEAD-DS. Each recording is 10s, single channel, 16-bit integers and sampled at 16 kHz.
Label Training Validation Testing All
cocktail_party 934 134 266 1334
interfering_speakers 733 105 209 1047
in_traffic 370 53 105 528
in_vehicle 409 59 116 584
music 1047 150 299 1496
quiet_indoors 368 53 104 525
reverberant_environment 156 22 44 222
wind_turbulence 307 44 88 439
speech_in_traffic 370 53 105 528
speech_in_vehicle 409 59 116 584
speech_in_music 1047 150 299 1496
speech_in_quiet_indoors 368 53 104 525
speech_in_reverberant_env 155 22 44 221
speech_in_wind_turbulence 307 44 88 439
Total 6980 1001 1987 9968
Table 4: AHEAD-DS uses the following data sources from CHiME 6 Dev [36] and HEAR-DS [15].
Source Sound Files Used / Total In Source
CHiME 6 Dev speech 44 / 52
HEAR-DS cocktail party 1334 / 1334
HEAR-DS in traffic 1056 / 1056
HEAR-DS in vehicle 1168 / 1168
HEAR-DS music 2992 / 2992
HEAR-DS quiet indoors 1050 / 1050
HEAR-DS reverb env 443 / 443
HEAR-DS wind turb 878 / 878
4.1 Data Processing Workflow
The data processing workflow is shown in Figure 1.
The sounds sourced from HEAR-DS and CHiME 6 Dev had been encoded in Waveform Audio File Format (WAV) as single-channel, 16-bit integers, and sampled at 16 kHz using several different equipment configurations as described above. AHEAD-DS retained this format.
The levels of CHiME 6 dev and music were high as they were optimised for playback. The signal levels of the other recordings were much lower, and not optimised for playback. The levels were standardised before mixing the clips. The audio files were converted into 32-bit floating-point format during standardisation and mixing, then converted back to 16-bit integers, and saved to disk. For everything except for CHiME 6 Dev and music, the audio was standardised by dividing by the RMS of cocktail party and multiplying by the RMS of CHiME 6 Dev. For music the audio was standardised by dividing by the RMS of music and multiplying by the RMS of CHiME 6 Dev.
These adjustments ensure that the levels of cocktail party and music were approximately similar to conversations in CHiME 6 Dev. Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev. Test results in a later section indicate that this procedure is critical to achieving good performance (See Gain Sensitivity Analysis section).
The first second from each recording in CHiME 6 dev is discarded, as it contains a loud beep to signal the start of the recording. The next 1100s were extracted from each file in 10s intervals. This generated 110 clips per file for a total of 4840 clips of speech.
For every environment from HEAR-DS, half the clips were mixed with speech, the other half are kept unmixed to represent the environment label. There are six environment labels which need to be mixed with speech. Of the 4840 speech clips taken from CHiME 6 Dev, 3793 clips were mixed with environmental sounds. 1047 clips were assigned to the label interfering speakers without further modifications.
For every five speech and environment sound file pairs, we mix them at 10,5,0,5,10-10,-5,0,5,10 dB signal-to-noise ratio (SNR) respectively. When speech and environment sounds were mixed, we boosted the levels of the quieter sound until it had equal RMS to the louder sound (this procedure occurs on top of standardisation). During mixing, speech was used as the signal and the environment sounds were used as noise. The speech levels were adjusted to the required SNR. The sounds of cocktail party and interfering speakers were used without mixing. No speech clip was used more than once.
The levels on the quieter sounds were boosted, as several sound clips for quiet indoors had very low RMS. When speech was mixed at 10-10 dB SNR, only near zero values were added. There were no or minimal signal changes after quantisation to 16-bit integers, when the mixed sound clips were saved back to disk. This causes sound clips with imperceptible speech being labelled with speech in quiet indoors, which is detrimental to model performance. The level boost ensures sound clips have acoustic realism and avoid labelling issues. Without the level boost, a human listener would not be able to determine there was speech in many speech in quiet indoors. With the level boost, background sounds are still nearly inaudible on manually listening, so the background acoustic signature is preserved.
After standardisation and mixing, we divide sounds for each label into training, validation, and testing sets using a percentage split of 70, 10, and 20 respectively.
Figure 1: A data flow diagram showing the data processing procedures. Half of the environment sounds were mixed with speech.
4.2 Data Augmentation
Data augmentation was applied as a regularisation technique [29]. It was applied on-the-fly only during training. Three augmentation procedures were applied, each independently with a probability of fifty percent: a gain of between -6 to 6 dB; noise to perturb the data, from a uniform distribution within range -0.003 to 0.003; and the length of the clip is shrunk or stretched between 0.9 and 1.1, using the Fourier resampler [25]. These augmentation procedures were chosen because they were fast to implement, and parameter ranges were small to perturb the data, without changing it so much that on manual listening it is unrecognisable from the original. In a later section we find augmentation improves performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
5 YAMNet+ Sound Recognition Model
The architecture diagram of YAMNet+ is shown in Figure 2. YAMNet+ is based on published YAMNet models [38] with the same internal architecture. We created a complete workflow for YAMNet+, as YAMNet was released with model definitions and weights but missing other components such as training code. We have released all code, making the implementation fully open source. For YAMNet+ we created training code, code to load weights, code for training via transfer learning, testing code, and conversion code to downsampling the model for deployment to edge computing platforms (We measured model speed using our test platform, a Google Pixel 3). YAMNet+ takes the waveform as input. A log mel spectrogram is computed from the waveform. The spectrogram is treated as a bitmap and processed using MobileNet [13]. The MobileNet starts with a convolutional block and a series of 13 blocks containing a depthwise separable convolution, batch normalisation, ReLU activation, convolution, batch normalisation and ReLU activation. There is a penultimate global average pooling layer. The final layer is a fully connected layer containing sigmoid activation of length equal to the number of class labels.
YAMNet+ slices the waveform into windows of length 960ms with a 480 millisecond overlap, therefore, a 10 second clip will be split into 20 windows. Zero padding is added to ensure the input waveform is of sufficient length. Sound recognition is performed on every 960 millisecond window; scores were not combined for a waveform divided across multiple windows; each window is treated separately. We use 960 milliseconds as it is a hyperparameter which is fixed by the use of transfer learning; this is essential to get good performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section). If a longer window is required, it is possible to work around the fixed length by adding the results from multiple windows and applying a softmax mathematical operation.
YAMNet was originally trained using AudioSet containing 521 class labels. Both YAMNet and YAMNet+ contain 3,751,369 parameters when the model is initialised using 521 class labels. There are 3,743,831 parameters when the model is initialised using 14 class labels, as used in HEAR-DS and AHEAD-DS.
Training occurred for a maximum of 100 epochs. An Adam optimiser [16] was used, with a learning rate of 0.00001. A decay-on-plateau learning rate scheduler was used. After 3 epochs, if performance on the validation set did not improve, the learning rate decays by multiplying by 0.5 after 6 epochs, if performance on the validation set did not improve, the training cycle was terminated early. The order of the training data was shuffled after every epoch.
Focal loss [20] with $\alpha=0.25$ and $\gamma=2$, was used. Focal loss down-weights easy and common classes, to better handle unbalanced datasets (AHEAD-DS is unbalanced).
A data loader was implemented to cycle through the data during training. It converts sound clips and labels into the format used during training. The sound clips are in WAV format and the labels are stored in a spreadsheet. Testing code was implemented to compute the confusion matrix, precision, recall, and mean average precision.
In order to optimise YAMNet+ for deployment on edge devices, the full-precision model was converted to a half-precision 16-bit floating-point model. This reduces the model size to 6.19 MiB. All subsequent testing was done using the 16-bit floating-point model.
Figure 2: The architecture of YAMNet+. The model computes the log mel spectrogram from the waveform. This is passed to a MobileNet which assigns confidence values to each label.
5.1 Label Smoothing
Label smoothing [22] with a smoothing factor of 0.1, was applied to the loss function. Label smoothing is a regularisation technique, which works by subtracting a small value from 1 for true classes, and adding a small value to 0 for incorrect classes. In a later section we show that label smoothing improves performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
5.2 Transfer Learning, Weight Initialisation, and Frozen Layers
YAMNet+ was initialised using transfer learning [41], where the weights were initially trained using AudioSet. Two ways of applying transfer learning were tested with YAMNet+. One method is to freeze all except the last layer, alternatively parameters in all layers are trainable. In a later section we will see having all layers trainable results in better performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
5.3 System Setup
YAMNet+ was developed using Python and TensorFlow. The development system was running Debian 12, FFmpeg 5.1.6, Python 3.11.2, TensorFlow 2.15.1 and 2.17.0, TensorFlow IO 0.37.1, and Keras 3. We make use of existing model definition code from YAMNet; this code is not fully compatible with newer versions of TensorFlow. We use the older TensorFlow to load and save a YAMNet model into a format which is compatible with newer TensorFlow. Then we use the newer TensorFlow to further tune this model.
TensorFlow Lite was used to convert the model to 16 bits to reduce the memory and computation requirements. The TensorFlow Lite model was deployed through an Android phone application [21]. The Android application was written using Flutter 3.24.5 and Dart 3.5.4.
The edge computing test system was a Google Pixel 3 phone running Android 13. The development system used a i7-1195G7 processor, 32 GiB of memory and a solid-state drive.
The models were trained using only the central processing unit (CPU). It was not possible to get graphics processing unit (GPU) acceleration due to compatibility issues with TensorFlow; however, GPU use would reduce training time significantly.
6 Test Results
Precision and recall curves are shown in Figure 3. The confusion matrix is shown in Figure 4. The metrics were produced using the test set of AHEAD-DS. Unless otherwise noted, tests were conducted using parameters outlined in the previous sections. Results shown in this section have been rounded to 2 decimal places.
A mean Average Precision (mAP) of 0.83 was achieved. This metric was calculated by averaging precision values across thresholds from 0.0 to 1.0 (in 0.1 increments) for each label, and then averaging those results. A threshold was applied to determine the label for each 960 millisecond window and compared against the ground truth. This gives the true positive, false positive, true negative, and false negative values. A higher mAP indicates better performance.
The confusion matrix indicates an accuracy of 0.93 was achieved. The confusion matrix was produced by computing the confidence of each label. The highest scoring label was compared to the ground truth. Each number in each cell represents the number of 960 millisecond windows assigned to a label. A higher number along the diagonals indicates better results.
A mAP of 0.83 and an accuracy of 0.93 indicates the model demonstrates very promising performance in scene recognition. There were a few labels which were misidentified a notable number of times. The labels speech in music and music were often misidentified during pauses in speech and only music could be heard. The same issue occurs for speech in traffic and in traffic. The label speech in vehicle was sometimes misidentified as speech in wind turbulence because the sound of rushing air entering the vehicle can often be heard.
Figure 3: YAMNet+ precision and recall curves. A mAP of 0.83 was achieved. A threshold was applied to determine the label for each 960 millisecond window and compared against the ground truth. A larger area under the curve indicates better performance.
Figure 4: YAMNet+ confusion matrix. An accuracy of 0.93 was achieved. The highest scoring label was compared to the ground truth. Each number in each cell represents the number of 960ms windows assigned to a label. A higher number along the diagonals indicates better results.
6.1 Learning Rate Sensitivity Analysis
The learning rate vs several validation metrics are shown in Figure 5. The best validation loss, mAP, and accuracy was achieved with a learning rate of 0.00001.
Learning rates of 0.001, 0.0001, 0.00001, and 0.000001 were validated. The previously described decay-on-plateau learning rate scheduler was used. A decay of 0.5 was applied after 3 epochs where validation does not improve, early termination after 6 epochs where validation does not improve. Tests were performed using transfer learning by initialsing with weights trained on AudioSet.
The following results were observed: validation losses of 1.0, 0.11, 0.11, and 0.12; mAP scores of 0.34, 0.82, 0.85, and 0.85; and accuracy values of 0.13, 0.93, 0.95, and 0.94. These outcomes correspond sequentially to learning rates of 0.001, 0.0001, 0.00001, and 0.000001.
The learning rate of 0.00001 emerged as the most favorable choice as it achieved the best overall results. The learning rate of 0.001 achieved the worst results; the resulting ML model was likely not practically useful. This was caused by the relatively large learning rate causing parameters to overshoot optimal values.
(a) Learning rate vs validation loss. Lower values are better. The Y-axis has a scale break, this prevents the substantial validation loss of 1.0 (at a learning rate of 0.001) from visually dominating the graph.
(b) Learning rate vs validation mAP. Higher values are better.
(c) Learning rate vs validation accuracy. Higher values are better.
Figure 5: The learning rate vs several metrics. The best validation loss, mAP, and accuracy was achieved with a learning rate of 0.00001.
6.2 Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation
The results of validating several configurations, by adjusting hyperparameters are shown in Figure 6. The best overall result occurs under the default configuration with validation loss, mAP, and accuracy of 0.11, 0.85, and 0.95 respectively. The default configuration had label smoothing enabled, augmentation enabled, all layers were trainable, and weight initialisation used weights trained on AudioSet instead of random initialisation. The learning rate was set to 0.00001 for all experiments in this section.
When label smoothing was not applied the ML model produced a validation loss, mAP, and accuracy of 0.02, 0.82, and 0.95 respectively (compared to 0.11, 0.85, and 0.95 when label smoothing was enabled). The configuration with no label smoothing produced a low validation loss of 0.02. However, this result did not lead to better mAP or accuracy; this occurs when predictions are wrong but still relatively close to the correct results. The application of label smoothing yielded improved overall results.
When augmentation was not applied the ML model produced validation loss, mAP, and accuracy of 0.12, 0.84, and 0.94 respectively (compared to 0.11, 0.85, and 0.95 when augmentation was enabled). Augmentation yielded a very small impact on performance.
When only the last layer remained unfrozen the ML model produced a validation loss, mAP, and accuracy of 0.18, 0.87, and 0.80 respectively (compared to 0.11, 0.85, and 0.95 when all layers were trainable). This configuration produced the highest mAP of 0.87. However, a high mAP can still be achieved even when there are many false negatives in the presence of few false positives. The default configuration had better validation loss and accuracy results. Therefore, having all layers trainable yielded improved overall results.
When using random weight initialisation, the ML model produced a validation loss, mAP, and accuracy of 0.53, 0.61, and 0.15 respectively (compared to 0.11, 0.85, and 0.95 when using weights trained on AudioSet). The metrics indicate the ML model produced with random weight initialisation was likely not practically useful. Initialising the weights, trained on AudioSet, through transfer learning was essential to achieving a performant ML model.
The implications of the small size of AHEAD-DS can be observed in the analysis of results using random weight initialisation. Random weight initialisation produced a high loss for both training and validation, 0.53 and 0.53 respectively. This indicates underfitting. The learning curves show that fine tuning weights trained on AudioSet is necessary, when faced with the limited training data in AHEAD-DS.
The tradeoff of using transfer learning is that several hyperparameters become fixed. Changing a hyperparameter impacts all parameters dependent on the said hyperparameter. A new model will need to be initialised and re-trained from scratch using AudioSet, before transfer learning can be applied. This makes it resource intensive to change hyperparameters such as extending the window length beyond 960ms.
(a) Configurations vs validation loss. Lower values are better. The configuration with no label smoothing produced the lowest validation loss. However, this result did not lead to better mAP or accuracy. This occurs when predictions are wrong but still relatively close to the correct results.
(b) Configurations vs validation mAP. Higher values are better. The configurations where only the last layer remained unfrozen produced the highest mAP. However, a high mAP can still be achieved even when there are many false negatives in the presence of few false positives. This configuration did not produce the highest accuracy and had high validation loss.
(c) Configurations vs validation accuracy. Higher values are better. The equal best accuracy, after rounding, was produced under the default configuration.
Figure 6: Configurations vs several validation metrics. The best overall result was produced under the default configuration with label smoothing enabled, augmentation enabled, all layers trainable, and weight initialisation with weights trained on AudioSet instead of random initialisation.
6.3 Gain Sensitivity Analysis
Tests were performed at different input gains. The validation mAP and accuracy are shown in Figure 7.
We validated with a gain of -20 to +20 dB in 5 dB increments. We observed best validation mAP and accuracy results when testing with 0 dB gain.Deviations in sound levels above or below the model’s anticipated range, as determined by its training data, led to diminished performance. This indicates adjustment of the levels is necessary to achieve optimal performance when applying the model to new datasets.
(a) Gain vs validation mAP. Higher values are better.
(b) Gain vs validation accuracy. Higher values are better.
Figure 7: The validation mAP and accuracy at different gain levels. The best results were achieved with zero gain. Higher mAP and accuracy indicate better results.
6.4 Model Speed and Edge Deployment
The mean processing times are shown in Figure 8. The tests were performed on a Google Pixel 3, a phone with modest specifications, released in 2018. We recorded 5, 10, 20, and 30s of audio on the phone, and measured the time it takes to process the audio by the model. For each length of time, we repeat the procedure five times and compute the mean time. We take a timestamp when the audio is buffered into memory and another timestamp when the results can be read from memory. The difference between the timestamps in milliseconds is recorded as the processing time. The processing time includes the time to compute the log mel spectrograms.
It takes a mean time of 200.4ms to process 5s of sound. There was approximately 50ms of latency to load the model, and approximately a linear increase of 30ms per 1 second of audio. In Figure 8, A straight line was fit to the data using linear regression. The line formula is $y=29.62x+48.31$. The Y-axis intercept is at 48.31. The line is almost a perfect fit to the data. This is a good indication that processing time increases linearly with recorded audio time. The latency of approximately 50ms is a one-off cost; so long as the variable representing the model remains in scope.
The tests were performed by deploying an Android application to the test Android phone. The application was launched and began buffering a maximum of 30s of audio. Audio from the beginning gets discarded as the most recent recording gets buffered. The user was then able to press a button and process a fixed interval of audio buffered in memory [21].
This processing time indicates that it is feasible to create an application which is able to perform scene recognition in real time, as it takes significantly less time to process the audio compared to the audio playback time. There are far newer and more performant edge devices compared to our test system, so we expect lower latency and processing time with newer devices.
Figure 8: The mean processing time for 5, 10, 20, and 30s of sound over five iterations on a Google Pixel 3. It takes a mean time of 200.4ms to process 5s of sound. There was approximately 50ms of latency to load the model, and a linear increase of 30ms per 1 second of audio. A straight line was fit to the data using linear regression. The line formula is $y=29.62x+48.31$. The Y-axis intercept is at 48.31. The line is almost a perfect fit to the data. This is a good indication that processing time increases linearly with recorded audio time.
7 Discussion
7.1 Summary of Findings
The experiments demonstrate that YAMNet+ achieves promising performance, with a mAP of 0.83 and an accuracy of 0.93. The study highlights two critical findings that are essential for YAMNet+ to perform well. One is the necessity of transfer learning from the pre-trained YAMNet model on AudioSet to achieve satisfactory performance. The other is the necessity of proper adjustment of sound levels to ensure optimal performance.
The real-time sound-based scene recognition capabilities of YAMNet+ on edge devices were successfully demonstrated. On a Google Pixel 3 smartphone (a phone with modest specifications, released in 2018), the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio. This indicates the feasibility of deploying such models for real-time applications on resource constrained devices.
The workflow and toolchain created for training and deploying a scene recognition model to edge devices is useful for future research in this area.
7.2 Comparison with Previous Work
Compared to the original YAMNet public release, YAMNet+ is an improvement because we implemented a complete workflow, so it can be tuned with transfer learning. YAMNet published without training code can only be used as a black box. YAMNet+ is also lightweight with 3.7 million parameters compared to the >80>80 million parameters of other sound recognition models [17] [10].
AHEAD-DS allows comparison of models using a standardised dataset with consistent labels, divided into training/validation/testing sets, and annotated with the ground truth. AHEAD-DS is particularly valuable given the difficulty of directly comparing existing studies, where datasets may be incomplete, not in a ready-to-use state, labels differ between datasets, and the code and datasets may not be publicly available.
7.3 Limitations
The testing performed in previous sections reveal several limitations of the YAMNet+ and AHEAD-DS. Both training and validation loss of 0.53 show the AHEAD-DS dataset is small in size, and leads to underfitting when training YAMNet+ from scratch using random weight initialisation. We worked around this problem by using transfer learning, where the model is initialised using weights trained on AudioSet.
Another implication of the small dataset is the observed small improvement after applying regularisation techniques such as label smoothing and augmentation. A model must generalise better to perform well on a larger dataset, therefore regularisation techniques should be more impactful if the AHEAD-DS is expanded.
The use of transfer learning comes with its own tradeoffs. We are locked to classifying 960 millisecond windows, as this is a hyperparameter which is fixed by the use of weights from an existing model. It is possible to work around the fixed length by adding the results from multiple windows and applying a softmax mathematical operation.
The tests of gain sensitivity analysis indicate adjustment of the levels is necessary to achieve optimal performance. Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev.
The final notable limitation is the dependency on TensorFlow. YAMNet+ is built using the TensorFlow ML library, as it is necessary to maintain compatibility with YAMNet weights, required for transfer learning. TensorFlow has GPU compatibility problems, disabling the use of a GPU for quick model training and complicating the codebase.
7.4 Next Steps
Future work can target three areas: the dataset, the model, and implementation details. The dataset is still relatively small compared to the likes of AudioSet and one finding is that transfer learning is essential to achieve reasonable performance. Therefore, the next step is to collect more data for a new version of AHEAD-DS. More sound clips, more varied sound environments, and more varied recording equipment configurations ensures models trained using a future version of AHEAD-DS are more generalisable.
We can improve the dataset by choosing scene labels in a systematic way. In this paper we only adopted labels from an existing study and looked at several studies which apply labels to sound clips. We found a lack of rigorous studies to explain why these labels are good choices. Examples such as speech in noise seem very generic and one talker on subway platform seem oddly specific. We have yet to determine the balance between labels being too specific or too generic, concerning the acoustic environments and listening contexts critical from an audiological perspective. This is important as good scene labels are crucial in hearing research, especially when conducting Ecological Momentary Assessment (EMA) [12]. EMAs are common in hearing research and involve surveying participants in real-time, and scene recognition models can automate this process, saving time and reducing the participants’ cognitive load. However, if these labels are ambiguous, inconsistent, or too broad/narrow, model accuracy will suffer, negating the benefits of automation and potentially increasing the burden on participants.
We may be able to improve the results by making better use of the limited data in AHEAD-DS. Speech and environment sounds can be left unmixed but standardised in the training set. More varied training data can be generated if they are mixed on-the-fly. So long as the test data is not mixed on-the-fly; results still remain comparable between test runs.
Things that would improve the model implementation would include rebuilding the model by first cleanly separating feature computation and model code. This will enable further experimentation with other features such as time domain based features. This may enable more effective detection of acoustic properties such as reverberations, which are more salient in the time domain. Rebuilding the model enables the ability to test different architectures such as EfficientNet or ResNet. Newer architectures such as EfficientNet might boost precision and recall.
We can improve the model implementation by using a modern ML library such as JAX or Pytorch. This allows us to avoid the complex legacy code and backward compatibility problems of TensorFlow. It would also significantly speed up training as more modern ML libraries have better compatibility with GPU acceleration.
We want to test the implementation on more types of edge hardware. This will allow us to determine the minimum resources necessary to enable real time sound recognition. If possible we want to deploy ML models to very power and resource constrained devices such as hearing devices and still maintain the ability to react to sounds in real time.
8 Conclusion
This study addresses key challenges in scene recognition for hearing devices by introducing a refined dataset and machine learning model. AHEAD-DS is a publicly available dataset specifically designed for audiologically relevant environments, offering a standardised dataset for comparing models with consistent labels and in a ready-to-use state. Complementing this, YAMNet+ is presented as a robust sound recognition model, optimised for deployment on resource-constrained edge devices like smartphones. YAMNet+ only requires 3.7 million parameters compared to >80>80 million parameters of other sound recognition models. It serves as an open-source baseline, including model definitions, training, and testing code. Benchmarking YAMNet+ on AHEAD-DS yielded a mAP of 0.83 and an accuracy of 0.93. The model’s efficiency was demonstrated on a Google Pixel 3 smartphone, processing each second of audio in approximately 30ms with 50ms latency. The established workflow and toolchain for training and deploying such models on edge devices offer a valuable resource for future advancements in this field.
9 Acknowledgements
We thank our colleagues at the Macquarie University Hearing Hub and Google Research for their invaluable help, and the Google Australia Catalyst Fund and Macquarie University for their support.
9.1 Funding
The authors disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was supported by the Google DFI Catalyst fund and Macquarie University.
9.2 Declaration of conflicting interests
The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
9.3 ORCID iDs
9.4 Links
Our website and code

### v4

@@
-A dataset and model for recognition of audiologically relevant environments for hearing aids: AHEAD-DS and YAMNet+
+A dataset and model for auditory scene recognition for hearing devices: AHEAD-DS and OpenYAMNet
 Abstract
-Scene recognition of audiologically relevant environments is important for hearing aids; however, it is challenging, in part because of the limitations of existing datasets.
+Scene recognition is important for hearing devices, however; this is challenging, in part because of the limitations of existing datasets.
 Datasets often lack public accessibility, completeness, or audiologically relevant labels, hindering systematic comparison of machine learning models.
-Deploying these models on resource-constrained edge devices presents another challenge.
-Our solution is two-fold: we leverage several open source datasets to create AHEAD-DS, a dataset designed for scene recognition of audiologically relevant environments, and introduce YAMNet+, a sound recognition model.
+Deploying such models on resource-constrained edge devices presents another challenge.The proposed solution is two-fold, a repack and refinement of several open source datasets to create AHEAD-DS, a dataset designed for auditory scene recognition for hearing devices, and introduce OpenYAMNet, a sound recognition model.
 AHEAD-DS aims to provide a standardised, publicly available dataset with consistent labels relevant to hearing aids, facilitating model comparison.
-YAMNet+ is designed for deployment on edge devices like smartphones connected to hearing devices, such as hearing aids and wireless earphones with hearing aid functionality; serving as a baseline model for sound-based scene recognition.
-YAMNet+ achieved a mean average precision of 0.83 and accuracy of 0.93 on the testing set of AHEAD-DS across fourteen categories of audiologically relevant environments.
-We found that applying transfer learning from the pretrained YAMNet model was essential.
-We demonstrated real-time sound-based scene recognition capabilities on edge devices by deploying YAMNet+ to an Android smartphone.
-Even with a Google Pixel 3 (a phone with modest specifications, released in 2018), the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio.
-Our website and code https://github.com/Australian-Future-Hearing-Initiative.
-Keywords: scene recognition, sound recognition, audiologically relevant labels, machine hearing, computer audition, AHEAD-DS, YAMNet+
-1 Introduction
-State of the art hearing devices incorporate scene recognition in order to identify audiological environments and apply appropriate processing (filters, amplification, beam forming, etc) to improve a user’s hearing experience [3].
+OpenYAMNet is designed for deployment on edge devices like smartphones connected to hearing devices, such as hearing aids and wireless earphones with hearing aid functionality, serving as a baseline model for sound-based scene recognition.
+OpenYAMNet achieved a mean average precision of 0.86 and accuracy of 0.93 on the testing set of AHEAD-DS across fourteen categories relevant to auditory scene recognition.
+Real-time sound-based scene recognition capabilities were demonstrated on edge devices by deploying OpenYAMNet to an Android smartphone.
+Even with a 2018 Google Pixel 3, a phone with modest specifications, the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio.
+The project website with links to code, data, and models.
+https://github.com/Australian-Future-Hearing-Initiative
+I Introduction
+State of the art hearing devices incorporate scene recognition in order to identify the auditory environment [34].
+Scene recognition can be used to trigger distinct hearing device processing strategies [3], such as wind reduction, beamforming, and music mode.
+Processing strategies improve a hearing device user’s experience by enhancing salient sounds while suppressing undesirable background noise.
 Several machine learning (ML) models have been proposed for scene recognition [23].
-However, there are several limitations: ML models are challenging to port to edge devices [4] and datasets have inconsistent labels (As outlined in Table 1).
-It is important that models can run on edge devices to preserve privacy by avoiding the transmission of sensitive data and to minimise the cost of data transmission.
-However, not all models are immediately deployable to devices, such as hearing aids and wireless earphones with hearing aid functionality (henceforth will be collectively referred to as hearing devices), or to portable devices such as smartphones that operate in tandem with hearing devices.
-Resource constraints and data format compatibility must be overcome for model deployment on the edge [23].
-Models are trained and tested using datasets with different scene labels and are thus not directly comparable.
-These datasets are either not public or incomplete, and general-purpose sound datasets do not have labels that are necessarily relevant to hearing devices [15].
-We were motivated to address these problems for several reasons: There is a gap in publicly available standardised datasets for scene recognition of audiologically relevant environments for hearing devices, and an open-source baseline ML model designed for edge devices is currently unavailable.
-Our ultimate goal is to improve the hearing experience for users of hearing devices.
-This paper makes the following contributions:
-  • A description of a sound dataset named Another HEaring AiD scenes Data Set (AHEAD-DS).
-This dataset was created for the purpose of scene recognition of audiologically relevant environments for hearing devices, which allows comparison of models using a standardised dataset with consistent labels.
-The dataset is released under a permissive licence, in a ready-to-use state, divided into training/validation/testing sets and annotated with the ground truth.
-  • A description of a ML model called YAMNet+ for sound based scene recognition.
+It is important that models can run locally to preserve privacy, by avoiding the transmission of sensitive data and to minimise the cost of data transmission.
+Hearing devices like hearing aids, wireless earphones with hearing aid functionality, and smartphones that operate in tandem with hearing devices are edge devices [8].
+Porting ML models to edge devices is challenging because such devices are resource constrained [4] and have limited data format compatibility [22].
+The challenge of scene recognition is further compounded by the gap in publicly available standardised audio scene recognition datasets.
+Existing datasets are either in-accessible, incomplete, or not annotated with labels relevant to hearing device listening situations.
+These limitations prevent the systematic evaluation of ML models for scene recognition.
+The motivation for this paper is to address the aforementioned challenges to improve scene recognition.
+In doing so, allow better application of processing strategies to improve the hearing experience for users of hearing devices.
+This paper makes the following three contributions.
+Firstly, a description of a sound dataset named Another HEaring AiD scenes Data Set (AHEAD-DS).
+This dataset was created for the purpose of scene recognition of the auditory environment, which allows comparison of models using a standardised dataset with consistent labels.
+The dataset is repacked and refined from several other open source datasets, released under a permissive licence, in a ready-to-use state, divided into training/validation/testing sets and annotated with the ground truth.
+Secondly, a description of a ML model called OpenYAMNet for sound based scene recognition.
 The model can act as a baseline model for further development of sound based scene recognition models.
 The model is released under a permissive licence, including model definitions, training, testing code, and code to convert it into a format deployable to edge devices.
-  • We implemented YAMNet+, trained it using AHEAD-DS, and presented benchmark results including precision, recall and inference speed.
-We have deployed the model to a smartphone to demonstrate it works on an edge device.
-This paper is organised as follows.
-The second section presents an overview of existing datasets and audiologically relevant environment labels.
-The third section presents an overview of several sound recognition models.
-The fourth section provides a description of AHEAD-DS and how the dataset was created.
-The fifth section describes the YAMNet+ sound recognition model.
-Section six presents the results of training and testing YAMNet+ using AHEAD-DS.
-Section seven discusses the overall findings of this paper and describes the possible future direction of the research.
-Section eight draws some conclusions.
-2 Scene Recognition Datasets for Hearing Devices
-There are several relevant studies (Table 1) involving scene recognition of audiologically relevant environments for hearing devices.
-The availability of a standardised dataset helped drive tremendous advancements in computer vision [2].
-It is hoped that a dataset for scene recognition for hearing devices will similarly drive advancements.
-Hence, we sought such a dataset, but noted the following challenges with existing datasets relevant to hearing devices.
-  • Datasets may be proprietary, with a restrictive licence and not uploaded to a publicly accessible repository. This was an issue with four studies [33, 8, 11, 39].
-  • Data used in the studies were not recorded. Data was processed on-the-fly and raw sound discarded in three studies [14, 18, 5].
-  • In one study [15], the dataset available online does not match the description in the published paper.
-  • The dataset is a mix of secondary sources, but the specific subset from these sources are unknown. This was an issue for three studies [15, 33, 31].
-  • The dataset is a mix of secondary sources but there is ambiguity in the mixing procedure.
-For example, there may be pseudo random procedures during processing of sound files, which are not clearly explained in the paper.
-This was an issue with two studies [15, 31].
-  • Each study uses datasets with different labels; therefore, they are not easily comparable. This is the case for all studies in Table 1.
-Additionally, several large general-purpose datasets exist for sound recognition.
+Finally, benchmark results of an implementation of the OpenYAMNet ML model, trained using AHEAD-DS, including precision, recall and inference speed.
+The ML model was deployed to a smartphone to demonstrate the model’s functionality on an edge device.
+The remainder of this paper is organised as follows.
+The remainder of this section presents related work through summaries of existing datasets and ML models.
+The section that follows presents the methods used to create AHEAD-DS and the OpenYAMNet model.
+The section after presents test results.
+The penultimate section discusses the findings of the test results, compares against existing work, and describes the possible future direction of the research.
+The final section draws some conclusions.
+I-A Existing Audio Scene Recognition Datasets for Hearing Devices
+Several existing datasets were collected as part of studies relevant to auditory scene recognition, a summary is shown in Table I. The table summarises advantages and issues for individual datasets.
+The main challenges of existing audio scene recognition datasets for hearing devices fall into three categories.
+Public accessibility, completeness, and labelling relevance.
+Accessibility was a problem for datasets which were immediately discarded after processing, or under a proprietary and restrictive licence.
+The issue of completeness stems from datasets being a mix of secondary sources, but the mixing procedure and specific subset from secondary sources are unknown.
+Several large general-purpose datasets exist for sound recognition.
 Datasets such as AudioSet [9], UrbanSound8K [28], ESC50 [26], and TAU scenes [36] provide extensive training material for ML models.
-However, labels are either too specific or too general to be used directly for the purpose here.
-For example, in AudioSet recordings can be labelled bus or truck.
-This level of granular sound recognition is unnecessary: it is only necessary to recognise and respond to motor vehicle noises to improve the hearing experience.
-However, data can still be used through transfer learning or labels can be remapped to more broad labels.
+However, the existing labels are either too specific or too general for the objective of auditory scene recognition for hearing devices.
+For example, AudioSet recordings can be labelled bus or truck.
+This level of granular sound recognition is unnecessary for a hearing device, it is only necessary to recognise and respond to broad categories like motor vehicle noises to improve the user’s experience.
+However, this data can still be used through transfer learning, or the labels can be remapped to a more general set.
 Conversely, TAU scenes provide location-based recordings which are too general, such as at airports and metro stations.
-The sound of a motor vehicle will be similarly disruptive irrespective of whether they occur at either location.
-However, labels could be modified and made more specific, but this would likely require listening to and re-annotating the recordings.
-UrbanSound8K and ESC50 provide extensive data.
-Both datasets contain very specific labels, such as the sounds from specific animals.
-Like AudioSet, this level of granularity is unnecessary for the purposes here.
-However, the data is still useful for transfer learning.
-Given these issues, we created our own data set using HEAR-DS as a starting point.
-Table 1: This table summarises relevant scene labels for hearing devices from several studies. Summarised are the data licence with notes on the data collection and processing procedures.
+The sound of a motor vehicle will be similarly disruptive irrespective of whether the sound occurs at either location.
+However, labels could be modified and made more specific, but the modification would likely require listening to and re-annotating the recordings.
+Another challenge is the use of different labels across datasets.
+It is difficult to directly compare ML models trained on each dataset and makes it difficult to combine the datasets into a single, more comprehensive one.
+TABLE I: This table summarises relevant datasets for auditory scene recognition for hearing devices.
 Study Labels Data Licence Notes
@@
 The classification system was a proprietary part of the hearing aids used in the study.
-[15] Cocktail party, Interfering speakers, In traffic, Speech in traffic, In vehicle, Speech in vehicle, Music, Speech in music, Quiet indoors, Speech in qu.
-Indoors, Reverberant env., Speech in reverb.
+[16] Cocktail party, Interfering speakers, In traffic, Speech in traffic, In vehicle, Speech in vehicle, Music, Speech in music, Quiet indoors, Speech in quiet indoors, Reverberant env., Speech in reverb.
 env.
@@
 YouTube source is proprietary.
+The data split and any preprocessing parameters are unknown.
 [7] Bus, Subway, Street, Indoor, Car Proprietary Sound dataset recorded by study authors, not released publicly.
@@
 [39] Speech in quiet, Speech in noise, Noise, Music Mix of open source and proprietary Sound dataset derived from proprietary Phonak Sound CD 2, self recordings on the London tube subway, and a song from royalty free freesound.org.
-3 Sound Recognition Machine Learning Models
-Three machine learning models relevant for sound recognition are presented in this section, PANN [17], AST [10], and YAMNet [38].
-The models in all of these studies were trained, validated, and tested using the AudioSet dataset.
-The findings are summarised in Table 2.
-The extensive training dataset makes any of the models suitable for tuning via transfer learning.
-While more sound recognition models exist, they are minor variations on the three aforementioned models [40].
-Most models make use of frequency domain based features, such as log mel spectrograms.
-They use either convolution or attention neural network blocks with dense layers.
-The spectrograms are treated as images and processed using an image processing neural network.
-Minor variations involve switching spectrograms with log mel spectrograms, or a different configuration of neural network blocks.
-A minority of models such as PANN use the raw waveform as a feature.
-The three models PANN, AST and YAMNet have the following tradeoffs.
-  • PANN and AST achieve better mean average precision (mAP) results compared to YAMNet, 0.44 and 0.46 vs 0.31 respectively, using the testing set of AudioSet.
-However, PANN and AST require over 20 times more parameters compared to YAMNet, approximately 81 and 86 million compared to 3.7 million respectively.
-This poses a challenge for deploying PANN and AST to resource-constrained edge devices.
-  • YAMNet is implemented using the TensorFlow ML framework, while PANN and AST are implemented using the Pytorch ML framework.
-A robust software pipeline exists to port any TensorFlow model to an edge device, such as an Android phone, using TensorFlow Lite.
-However, YAMNet’s implementation is not fully open source, while model definitions and weights have been released, training code is missing from the code repositories.
-To address these issues, we developed our own model by creating a variation of YAMNet.
-Table 2: This table details several machine learning models for sound recognition.
-The mAP for the AudioSet testing set and number of parameters are those reported by the authors of each study.
-PANN and AST have higher mAP and use significantly more parameters compared to YAMNet/YAMNet+.
-This suggests that while PANN and AST have better sound recognition capabilities, they demand significantly more computational resources.
-Model Features Architecture Parameters mAP
-YAMNet [38] / YAMNet+ Log mel spectrogram Convolution 3,751,369 0.31
-PANN [17] Log mel spectrogram, waveform Convolution 81,065,487 0.44
-[…]
-4 Another HEaring AiD scenes Data Set (AHEAD-DS)
-We leverage several open source datasets to create a dataset which we called AHEAD-DS.
-A summary of the dataset is shown in Table 3.
-A summary of the sources of the dataset is shown in Table 4.
+[36] Airport, Bus, Metro, Metro station, Park, Public square, Shopping mall, Street pedestrian, Street traffic, Tram Open source The TAU urban audio visual scenes datasets were recorded in several European cities.”
+I-B Existing Sound Recognition Models
+Several existing sound recognition models exist, the three presented in this section are Yet Another Mobile Network (YAMNet) [27], Perch [10], and Audio Spectrogram Transformer (AST) [11].
+The three representative models were selected as the models provide a good cross-section of common deep learning designs, many other ML models are variations of the three architectures [40].
+AST and YAMNet were pre-trained, validated, and tested using the AudioSet dataset, while Perch was pre-trained using XenoCanto [33].
+The extensive datasets used for pre-training makes any of the models suitable for fine tuning via transfer learning.
+AST and YAMNet use log mel spectrograms as features, while Perch uses a per-channel energy normalised mel spectrogram.
+The spectrograms are represented as two-dimensional inputs and processed using neural networks designed for image recognition.
+YAMNet’s implementation is light weight at 3.7 million parameters compared to Perch and AST, which use 7.8 and 86 million parameters respectively.
+This makes YAMNet a good match for edge computing.
+However, YAMNet’s implementation is open weight but not fully open source, since only model definitions and weights have been released.
+The training code for YAMNet is missing from the code repositories, while Perch and AST are fully open source.
+II Methods
+This section outlines the methodology used to create the AHEAD-DS dataset and the OpenYAMNet model.
+The procedures for data curation, preprocessing, and labelling for AHEAD-DS are detailed, followed by a description of the OpenYAMNet model architecture, training, and deployment on edge devices.
+II-A Another HEaring AiD scenes Data Set (AHEAD-DS)
+AHEAD-DS repackages and refines the open source datasets HEAR-DS and CHiME 6 Dev [37].
+A summary of the dataset is shown in Table II.
 There are a total of 9968 recordings, 10s per recording, across 14 different labels.
 Each instance is single-channel, 16-bit integers and sampled at 16 kHz.
-Despite the limitations outlined earlier, AHEAD-DS is based on HEAR-DS and CHiME 6 Dev [36]. HEAR-DS provides a solid foundation for the following reasons:
-  • The recordings of environment sounds have been open sourced with a permissive licence.
-  • According to the authors of HEAR-DS, the 14 labels have been chosen from audiologically relevant environments for hearing devices by audiologists and Audifon, a manufacturer of hearing aids.
-AHEAD-DS follows the design of HEAR-DS, it employs an identical set of 14 labels.
+AHEAD-DS follows the design of HEAR-DS and employs an identical set of 14 labels.
 This set comprises six environmental sound categories and an additional six categories representing speech instances within these respective environments.
-Hearing devices likely require different signal processing strategies to optimise speech intelligibility in each acoustic setting [3].
-Therefore, the detection of speech embedded within specific environments represents a step towards creating an improved hearing experience for a hearing devices user.
-The main issue with HEAR-DS is that the dataset is not in a fixed, ready-to-use state.
-This is because the published data is unmixed, and some pseudo-random procedures were used during the production of the HEAR-DS dataset.
-Since the code has not been published, these procedures cannot be fully reproduced.
-This causes the following problems.
-  • It is not possible to determine the split for training and testing.
-  • It is not possible to produce results which can be compared directly against the ones in the HEAR-DS paper.
-The advantage of AHEAD-DS is that the dataset is in a ready-to-use state.
-The code for the mixing procedure is published and all mixing parameters are documented and reproducible.
-This is important because it removes a variable when comparing ML models.
-The speech and environment sounds have been released in both mixed and unmixed forms, eliminating any ambiguity regarding the subsets used to derive our new dataset.
 The sounds from each environment for AHEAD-DS were taken from HEAR-DS.
@@
 The head was equipped with a synthetic ear, the GRAS KB1065/1066 Pinnae.
-An Audiofon microphone was placed in each ear of the dummy head in an in-the-canal position.
+An Audifon microphone was placed in each ear of the dummy head in an in-the-canal position.
 A Focusrite Scarlett 18i6 sound interface was used for analogue to digital conversion.
@@
 There were a total of 2992 clips of 10s of music.
-One issue raised about AHEAD-DS and HEAR-DS was the constant acoustic signature of the reverberant environment might interfere with the classification of the acoustic signature of the other environments.
-For sound recognition, if a human can identify the presence of a reverberant environment, an acoustic signature exists for reverberant environments.
-Deep learning will automatically generate compensating features; these features are issues of concern for manual feature engineering; but not for deep learning [19].
 CHiME 6 Dev was used for speech samples in AHEAD-DS instead of CHiME 5, which was used in HEAR-DS.
 Both contain the same speech recordings, but the start and end times of all recordings have been corrected in CHiME 6 Dev to ensure all recordings are synchronised.
-CHiME 6 dev is a subset of CHiME 6; it is smaller than the whole dataset but sufficiently long for mixing with environment sounds to provide all the speech in $\langle environment\rangle$ clips and to produce interfering speakers.
+CHiME 6 dev is a subset of CHiME 6, it is smaller than the whole dataset but sufficiently long for mixing with environment sounds to provide all the speech in $\langle environment\rangle$ clips and to produce interfering speakers.
 The interfering speakers clips were derived entirely from CHiME 6 dev.
@@
 A total 44 sound files from CHiME 6 Dev were used in AHEAD-DS, 24 recordings from the first environment and 20 in the second.
-Each session consisted of four people and would begin by having participants take turns to introduce themselves via speaking and then read out a short passage.
-Once this is complete, participants engage in casual conversation as they prepare and share a meal together.
+Each session consisted of four people and would begin by having participants take turns to introduce the participants via speaking and then read out a short passage.
+Once the introduction is complete, participants engage in casual conversation as the participants prepare and share a meal together.
 The authors of CHiME 6 dev made the recordings by placing six Microsoft Kinects throughout the home in the first session and five in the second.
@@
 In CHiME 6 additional speech was recorded using body-worn microphones, but only the recordings from the Kinect were used.
-The recordings from body-worn microphones lacked acoustic realism as they sounded like clean studio recordings.
+The recordings from body-worn microphones lacked acoustic realism as the recordings sounded like clean studio recordings.
 The label interfering speakers contains only speech taken from CHiME 6 and is not mixed with environment sounds.
-It represents a scenario with multiple speakers where the speech is intelligible.
+The label represents a scenario with multiple speakers where the speech is intelligible.
 The label cocktail party contains multiple speakers but the speech is unintelligible.
-Compared to datasets like AudioSet, AHEAD-DS is still relatively small, which is a disadvantage.
-In later sections, we analyse the implications of having a small dataset (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
-Table 3: This table summarises the contents of AHEAD-DS. Each recording is 10s, single channel, 16-bit integers and sampled at 16 kHz.
+TABLE II: This table summarises the contents of AHEAD-DS. Each recording is 10s, single channel, 16-bit integers and sampled at 16 kHz.
 Label Training Validation Testing All
@@
 speech_in_quiet_indoors 368 53 104 525
-speech_in_reverberant_env 155 22 44 221
+speech_in_reverb_env 155 22 44 221
 speech_in_wind_turbulence 307 44 88 439
 Total 6980 1001 1987 9968
-Table 4: AHEAD-DS uses the following data sources from CHiME 6 Dev [36] and HEAR-DS [15].
-Source Sound Files Used / Total In Source
-CHiME 6 Dev speech 44 / 52
-HEAR-DS cocktail party 1334 / 1334
-HEAR-DS in traffic 1056 / 1056
-HEAR-DS in vehicle 1168 / 1168
-HEAR-DS music 2992 / 2992
-HEAR-DS quiet indoors 1050 / 1050
-HEAR-DS reverb env 443 / 443
-HEAR-DS wind turb 878 / 878
-4.1 Data Processing Workflow
-The data processing workflow is shown in Figure 1.
-The sounds sourced from HEAR-DS and CHiME 6 Dev had been encoded in Waveform Audio File Format (WAV) as single-channel, 16-bit integers, and sampled at 16 kHz using several different equipment configurations as described above.
+II-B AHEAD-DS Data Processing Workflow
+The workflow to create AHEAD-DS is shown in Figure 1.
+The sounds sourced from HEAR-DS and CHiME 6 Dev had been encoded in Waveform Audio File Format (WAV) as single-channel, 16-bit integers, and sampled at 16 kHz using several different equipment configurations.
 AHEAD-DS retained this format.
-The levels of CHiME 6 dev and music were high as they were optimised for playback.
+For all level adjustments and mixing, the audio files were converted into 32-bit floating-point format during processing.
+The values were then clipped to the values of $int16_{max}$ and $int16_{min}$ to avoid quantisation errors, then converted back to 16-bit integers, before being saved to disk.
+The levels of CHiME 6 dev and music were high as the levels were optimised for playback.
 The signal levels of the other recordings were much lower, and not optimised for playback.
 The levels were standardised before mixing the clips.
-The audio files were converted into 32-bit floating-point format during standardisation and mixing, then converted back to 16-bit integers, and saved to disk.
-For everything except for CHiME 6 Dev and music, the audio was standardised by dividing by the RMS of cocktail party and multiplying by the RMS of CHiME 6 Dev.
+For everything except CHiME 6 Dev and music, the audio was standardised by dividing by the RMS of cocktail party and multiplying by the RMS of CHiME 6 Dev.
 For music the audio was standardised by dividing by the RMS of music and multiplying by the RMS of CHiME 6 Dev.
@@
 Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev.
-Test results in a later section indicate that this procedure is critical to achieving good performance (See Gain Sensitivity Analysis section).
-The first second from each recording in CHiME 6 dev is discarded, as it contains a loud beep to signal the start of the recording.
+The first second from each recording in CHiME 6 dev is discarded, as the first second contains a loud beep to signal the start of the recording.
 The next 1100s were extracted from each file in 10s intervals.
@@
 1047 clips were assigned to the label interfering speakers without further modifications.
-For every five speech and environment sound file pairs, we mix them at 10,5,0,5,10-10,-5,0,5,10 dB signal-to-noise ratio (SNR) respectively.
-When speech and environment sounds were mixed, we boosted the levels of the quieter sound until it had equal RMS to the louder sound (this procedure occurs on top of standardisation).
+Every five speech and environment sound file pairs are mixed at 10,5,0,5,10-10,-5,0,5,10 dB signal-to-noise ratio (SNR).
+The mixing function is shown in Algorithm 1.
+The reason to mix speech at 10-10 dB SNR was to allow for the possibility to detect speech, though not intelligibility, even in edge cases that would challenge a human with good hearing.
+When speech and environment sounds were mixed, the levels of the quieter sound were boosted until it had equal RMS to the louder sound, this procedure occurs on top of standardisation.
 During mixing, speech was used as the signal and the environment sounds were used as noise.
@@
 No speech clip was used more than once.
-The levels on the quieter sounds were boosted, as several sound clips for quiet indoors had very low RMS.
-When speech was mixed at 10-10 dB SNR, only near zero values were added.
-There were no or minimal signal changes after quantisation to 16-bit integers, when the mixed sound clips were saved back to disk.
-This causes sound clips with imperceptible speech being labelled with speech in quiet indoors, which is detrimental to model performance.
-The level boost ensures sound clips have acoustic realism and avoid labelling issues.
-Without the level boost, a human listener would not be able to determine there was speech in many speech in quiet indoors.
-With the level boost, background sounds are still nearly inaudible on manually listening, so the background acoustic signature is preserved.
-After standardisation and mixing, we divide sounds for each label into training, validation, and testing sets using a percentage split of 70, 10, and 20 respectively.
-Figure 1: A data flow diagram showing the data processing procedures. Half of the environment sounds were mixed with speech.
-4.2 Data Augmentation
-Data augmentation was applied as a regularisation technique [29].
-It was applied on-the-fly only during training.
-Three augmentation procedures were applied, each independently with a probability of fifty percent: a gain of between -6 to 6 dB; noise to perturb the data, from a uniform distribution within range -0.003 to 0.003; and the length of the clip is shrunk or stretched between 0.9 and 1.1, using the Fourier resampler [25].
-These augmentation procedures were chosen because they were fast to implement, and parameter ranges were small to perturb the data, without changing it so much that on manual listening it is unrecognisable from the original.
-In a later section we find augmentation improves performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
-5 YAMNet+ Sound Recognition Model
-The architecture diagram of YAMNet+ is shown in Figure 2.
-YAMNet+ is based on published YAMNet models [38] with the same internal architecture.
-We created a complete workflow for YAMNet+, as YAMNet was released with model definitions and weights but missing other components such as training code.
-We have released all code, making the implementation fully open source.
-For YAMNet+ we created training code, code to load weights, code for training via transfer learning, testing code, and conversion code to downsampling the model for deployment to edge computing platforms (We measured model speed using our test platform, a Google Pixel 3).
-YAMNet+ takes the waveform as input.
+The reason to boost the quieter of speech or environment sound is to avoid the following situation.
+When speech was mixed at 10-10 dB SNR with quiet environment sounds such as quiet indoors, speech was imperceptible.
+This caused sound clips with imperceptible speech being labelled with a speech in environment label e.g.
+speech in quiet indoors, which is detrimental to model performance.
+The level boost ensures sound clips match the label.
+After standardisation and mixing, sounds are divided into training, validation, and testing sets using a percentage split of 70, 10, and 20 respectively.
+Figure 1: A data flow diagram showing the data processing workflow.
+Algorithm 1 Pseudocode to mix speech with environment sound.
+1:function Mix($speech,env,snr$)
+2:  $speech_{rms}\leftarrow\text{RMS}(speech)$
+3:  $env_{rms}\leftarrow\text{RMS}(env)$
+4:  $\triangleright$ Prevent division by zero
+5:  if $speech_{rms}==0$ then
+6:   $speech_{rms}\leftarrow 1e-6$
+7:  end if
+8:  if $env_{rms}==0$ then
+9:   $env_{rms}\leftarrow 1e-6$
+10:  end if
+11:  $\triangleright$ Normalise amplitudes
+12:  if $speech_{rms}>env_{rms}$ then
+13:   $speech_{boosted}\leftarrow speech$
+14:   $env_{boosted}\leftarrow env\times\frac{speech_{rms}}{env_{rms}}$
+15:  else
+16:   $speech_{boosted}\leftarrow speech\times\frac{env_{rms}}{speech_{rms}}$
+17:   $env_{boosted}\leftarrow env$
+18:  end if
+19:  $\triangleright$ Calculate gain from SNR
+20:  $gain_{factor}\leftarrow 10^{\left(\frac{snr}{20}\right)}$
+21:  $\triangleright$ Apply gain and mix
+22:  $speech_{gained}\leftarrow speech_{boosted}\times gain_{factor}$
+23:  $mixed\leftarrow speech_{gained}+env_{boosted}$
+24:  return mixedmixed
+25:end function
+II-C OpenYAMNet Model
+The architecture diagram of OpenYAMNet is shown in Figure 2.
+OpenYAMNet is based on YAMNet [27] with the same internal architecture.
+A complete workflow for OpenYAMNet was created, as the original YAMNet was released with model definitions and weights but missing other components such as training code.
+All code has been released, making the implementation fully open source.
+For OpenYAMNet the code consists of training code, code to load weights, code for training via transfer learning, testing code, and conversion code to downsampling the model for deployment to edge computing platforms.
+OpenYAMNet takes the waveform as input.
 A log mel spectrogram is computed from the waveform.
-The spectrogram is treated as a bitmap and processed using MobileNet [13].
-The MobileNet starts with a convolutional block and a series of 13 blocks containing a depthwise separable convolution, batch normalisation, ReLU activation, convolution, batch normalisation and ReLU activation.
+The spectrogram is treated as two dimension input and processed using MobileNet [14].
+MobileNet starts with a convolutional block and a series of 13 blocks containing a depthwise separable convolution, batch normalisation, ReLU activation, convolution, batch normalisation and ReLU activation.
 There is a penultimate global average pooling layer.
 The final layer is a fully connected layer containing sigmoid activation of length equal to the number of class labels.
-YAMNet+ slices the waveform into windows of length 960ms with a 480 millisecond overlap, therefore, a 10 second clip will be split into 20 windows.
+OpenYAMNet slices the waveform into windows of length 960ms with a 480 millisecond overlap, therefore, a 10 second clip will be split into 20 windows.
 Zero padding is added to ensure the input waveform is of sufficient length.
-Sound recognition is performed on every 960 millisecond window; scores were not combined for a waveform divided across multiple windows; each window is treated separately.
-We use 960 milliseconds as it is a hyperparameter which is fixed by the use of transfer learning; this is essential to get good performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
-If a longer window is required, it is possible to work around the fixed length by adding the results from multiple windows and applying a softmax mathematical operation.
-YAMNet was originally trained using AudioSet containing 521 class labels.
-Both YAMNet and YAMNet+ contain 3,751,369 parameters when the model is initialised using 521 class labels.
-There are 3,743,831 parameters when the model is initialised using 14 class labels, as used in HEAR-DS and AHEAD-DS.
+Scene recognition is performed on every 960 millisecond window and each window is treated separately.
+The 960 millisecond window length is a hyperparameter which is fixed by the use of transfer learning.
+Figure 2: The architecture of OpenYAMNet.
+II-D OpenYAMNet Training Workflow
+OpenYAMNet weights were initialised using transfer learning [41], taken from YAMNet trained on AudioSet.
 Training occurred for a maximum of 100 epochs.
-An Adam optimiser [16] was used, with a learning rate of 0.00001.
+Focal loss [19] with $\alpha=0.25$ and $\gamma=2$ was used as the loss function.
+An Adam optimiser [17] was used, with an initial learning rate of 0.00001.
 A decay-on-plateau learning rate scheduler was used.
-After 3 epochs, if performance on the validation set did not improve, the learning rate decays by multiplying by 0.5 after 6 epochs, if performance on the validation set did not improve, the training cycle was terminated early.
-The order of the training data was shuffled after every epoch.
-Focal loss [20] with $\alpha=0.25$ and $\gamma=2$, was used. Focal loss down-weights easy and common classes, to better handle unbalanced datasets (AHEAD-DS is unbalanced).
-A data loader was implemented to cycle through the data during training.
-It converts sound clips and labels into the format used during training.
-The sound clips are in WAV format and the labels are stored in a spreadsheet.
-Testing code was implemented to compute the confusion matrix, precision, recall, and mean average precision.
-In order to optimise YAMNet+ for deployment on edge devices, the full-precision model was converted to a half-precision 16-bit floating-point model.
-This reduces the model size to 6.19 MiB.
-All subsequent testing was done using the 16-bit floating-point model.
-Figure 2: The architecture of YAMNet+. The model computes the log mel spectrogram from the waveform. This is passed to a MobileNet which assigns confidence values to each label.
-5.1 Label Smoothing
-Label smoothing [22] with a smoothing factor of 0.1, was applied to the loss function.
-Label smoothing is a regularisation technique, which works by subtracting a small value from 1 for true classes, and adding a small value to 0 for incorrect classes.
-In a later section we show that label smoothing improves performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
-5.2 Transfer Learning, Weight Initialisation, and Frozen Layers
-YAMNet+ was initialised using transfer learning [41], where the weights were initially trained using AudioSet.
-Two ways of applying transfer learning were tested with YAMNet+.
-One method is to freeze all except the last layer, alternatively parameters in all layers are trainable.
-In a later section we will see having all layers trainable results in better performance (See Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation section).
-5.3 System Setup
-YAMNet+ was developed using Python and TensorFlow.
-The development system was running Debian 12, FFmpeg 5.1.6, Python 3.11.2, TensorFlow 2.15.1 and 2.17.0, TensorFlow IO 0.37.1, and Keras 3.
-We make use of existing model definition code from YAMNet; this code is not fully compatible with newer versions of TensorFlow.
-We use the older TensorFlow to load and save a YAMNet model into a format which is compatible with newer TensorFlow.
-Then we use the newer TensorFlow to further tune this model.
-TensorFlow Lite was used to convert the model to 16 bits to reduce the memory and computation requirements.
-The TensorFlow Lite model was deployed through an Android phone application [21].
-The Android application was written using Flutter 3.24.5 and Dart 3.5.4.
-The edge computing test system was a Google Pixel 3 phone running Android 13. The development system used a i7-1195G7 processor, 32 GiB of memory and a solid-state drive.
+After 3 epochs, if performance on the validation set did not improve, the learning rate decays by multiplying by 0.5.
+After 6 epochs, if performance on the validation set did not improve, the training cycle was terminated.
+The order of the training data was shuffled at the end of every epoch.
+Label smoothing [21] with a smoothing factor of 0.1 was applied as a regularisation function.
+Data augmentation [29] was also applied as a regularisation technique.
+Three augmentation procedures were applied, each independently with a probability of fifty percent: a gain of between -6 to 6 dB, noise to perturb the data, from a uniform distribution within range -0.003 to 0.003, and the length of the clip is shrunk or stretched between 0.9 and 1.1, using the Fourier resampler [24].
+The window length of 960 milliseconds and overlap of 480 milliseconds causes each 10 second audio clip in AHEAD-DS to be divided into 20 windows.
+The dataset contains 6,980 training, 1,001 validation, and 1,987 testing clips.
+A total of 139,600 windows are used for training, 20,020 for validation, and 39,740 for testing.
+II-E Software Environment
+OpenYAMNet was developed using Python and TensorFlow.
+For technical details, links to the code are provided in the Data availability statement.
 The models were trained using only the central processing unit (CPU).
-It was not possible to get graphics processing unit (GPU) acceleration due to compatibility issues with TensorFlow; however, GPU use would reduce training time significantly.
-6 Test Results
+It was not possible to get graphics processing unit (GPU) acceleration due to compatibility issues with TensorFlow, however, GPU use would reduce training time significantly.
+In order to optimise OpenYAMNet for deployment on edge devices, the full-precision model was converted to a half-precision 16-bit floating-point model.
+This was done using TensorFlow Lite.
+The model was deployed through a phone application [20] written using Flutter.
+The edge computing test system was a Google Pixel 3 running Android.
+III Results
+This section outlines the results of testing OpenYAMNet using AHEAD-DS.
+The analysis includes the model’s performance on precision, recall, and accuracy.
+It also compares the impact of various hyperparameters, evaluates the processing time on an edge platform, and benchmarks the model against several other ML models.
+III-A Precision, Recall, and Accuracy
 Precision and recall curves are shown in Figure 3.
@@
 Results shown in this section have been rounded to 2 decimal places.
-A mean Average Precision (mAP) of 0.83 was achieved.
-This metric was calculated by averaging precision values across thresholds from 0.0 to 1.0 (in 0.1 increments) for each label, and then averaging those results.
-A threshold was applied to determine the label for each 960 millisecond window and compared against the ground truth.
+A mAP of 0.86 was achieved.
+All mAP metrics in this paper were calculated by averaging precision values across thresholds from 0.0 to 1.0, in 0.01 increments and then averaging those over all labels.
 This gives the true positive, false positive, true negative, and false negative values.
@@
 The highest scoring label was compared to the ground truth.
-Each number in each cell represents the number of 960 millisecond windows assigned to a label.
 A higher number along the diagonals indicates better results.
-A mAP of 0.83 and an accuracy of 0.93 indicates the model demonstrates very promising performance in scene recognition.
-There were a few labels which were misidentified a notable number of times.
-The labels speech in music and music were often misidentified during pauses in speech and only music could be heard.
-The same issue occurs for speech in traffic and in traffic.
-The label speech in vehicle was sometimes misidentified as speech in wind turbulence because the sound of rushing air entering the vehicle can often be heard.
-Figure 3: YAMNet+ precision and recall curves.
-A mAP of 0.83 was achieved.
-A threshold was applied to determine the label for each 960 millisecond window and compared against the ground truth.
+(a) OpenYAMNet precision and recall curves for cocktail party, interfering speaker, and several environments.
+(b) OpenYAMNet precision and recall curves for speech in several environments.
+Figure 3: OpenYAMNet precision and recall curves.
+A mAP of 0.86 was achieved on the test set.
+Each 960 millisecond window contributed a prediction outcome.
 A larger area under the curve indicates better performance.
-Figure 4: YAMNet+ confusion matrix.
-An accuracy of 0.93 was achieved.
-The highest scoring label was compared to the ground truth.
-Each number in each cell represents the number of 960ms windows assigned to a label.
+Figure 4: OpenYAMNet confusion matrix.
+An accuracy of 0.93 was achieved on the test set.
+Each 960ms window contributed a prediction outcome.
 A higher number along the diagonals indicates better results.
-6.1 Learning Rate Sensitivity Analysis
-The learning rate vs several validation metrics are shown in Figure 5. The best validation loss, mAP, and accuracy was achieved with a learning rate of 0.00001.
-Learning rates of 0.001, 0.0001, 0.00001, and 0.000001 were validated.
-The previously described decay-on-plateau learning rate scheduler was used.
-A decay of 0.5 was applied after 3 epochs where validation does not improve, early termination after 6 epochs where validation does not improve.
-Tests were performed using transfer learning by initialsing with weights trained on AudioSet.
-The following results were observed: validation losses of 1.0, 0.11, 0.11, and 0.12; mAP scores of 0.34, 0.82, 0.85, and 0.85; and accuracy values of 0.13, 0.93, 0.95, and 0.94.
-These outcomes correspond sequentially to learning rates of 0.001, 0.0001, 0.00001, and 0.000001.
-The learning rate of 0.00001 emerged as the most favorable choice as it achieved the best overall results.
-The learning rate of 0.001 achieved the worst results; the resulting ML model was likely not practically useful.
-This was caused by the relatively large learning rate causing parameters to overshoot optimal values.
-(a) Learning rate vs validation loss.
-Lower values are better.
-The Y-axis has a scale break, this prevents the substantial validation loss of 1.0 (at a learning rate of 0.001) from visually dominating the graph.
-(b) Learning rate vs validation mAP. Higher values are better.
-(c) Learning rate vs validation accuracy. Higher values are better.
-Figure 5: The learning rate vs several metrics. The best validation loss, mAP, and accuracy was achieved with a learning rate of 0.00001.
-6.2 Label Smoothing, Augmentation, Unfreezing Only the Last Layer, and Random Weight Initialisation
-The results of validating several configurations, by adjusting hyperparameters are shown in Figure 6.
-The best overall result occurs under the default configuration with validation loss, mAP, and accuracy of 0.11, 0.85, and 0.95 respectively.
-The default configuration had label smoothing enabled, augmentation enabled, all layers were trainable, and weight initialisation used weights trained on AudioSet instead of random initialisation.
+III-B Learning Rate Sensitivity Analysis
+To evaluate learning rate sensitivity, four values were validated: 0.001, 0.0001, 0.00001, and 0.000001.
+For each learning rate, models were trained three times with different random seeds to obtain the mean and standard deviation (SD) for key metrics: loss, mAP, and accuracy on the validation set.
+The previously described decay-on-plateau learning rate scheduler was used, and model weights were initialised via transfer learning with weights pre-trained on the AudioSet dataset.
+The results showed that learning rates of 0.0001, 0.00001, and 0.000001 all produced very similar performance, with a mean loss of 0.11, mean mAP ranging from 0.85 to 0.88, and a mean accuracy of 0.94.
+The low SD, which rounded to zero for several metrics, highlighted the consistency of these results.
+The learning rate of 0.001 performed poorly with a mean validation loss of 2.0, mean validation mAP of 0.5, and Mean Validation Accuracy of 0.14.
+The geometric mean of the three well performing learning rates, 0.00001 was chosen as the default learning rate.
+TABLE III: The mean and standard deviations for loss, mAP, and accuracy over three iterations of each learning rate on the validation set.
+Learning Rate Mean Validation Loss $\pm$ SD Mean Validation mAP $\pm$ SD Mean Validation Accuracy $\pm$ SD
+0.001 2.00 $\pm$ 1.09 0.50 $\pm$ 0.27 0.14 $\pm$ 0.01
+0.0001 0.11 $\pm$ 0.00 0.85 $\pm$ 0.01 0.94 $\pm$ 0.00
+0.00001 0.11 $\pm$ 0.00 0.88 $\pm$ 0.00 0.94 $\pm$ 0.01
+0.000001 0.11 $\pm$ 0.00 0.88 $\pm$ 0.00 0.94 $\pm$ 0.00
+III-C Ablation Study
+The results of an ablation study validating several configurations are shown in Table IV.
+Each metric was calculated by running 10 iterations using bootstrapping, where 200 samples were chosen with replacement from the validation set.
+The result under the default configuration achieved a mean validation loss, mAP, and accuracy of 0.11, 0.88, and 0.95 respectively.
+The default configuration had label smoothing enabled, augmentation enabled, all layers were trainable, and weight initialisation used weights pre-trained on AudioSet.
 The learning rate was set to 0.00001 for all experiments in this section.
-When label smoothing was not applied the ML model produced a validation loss, mAP, and accuracy of 0.02, 0.82, and 0.95 respectively (compared to 0.11, 0.85, and 0.95 when label smoothing was enabled).
-The configuration with no label smoothing produced a low validation loss of 0.02.
-However, this result did not lead to better mAP or accuracy; this occurs when predictions are wrong but still relatively close to the correct results.
-The application of label smoothing yielded improved overall results.
-When augmentation was not applied the ML model produced validation loss, mAP, and accuracy of 0.12, 0.84, and 0.94 respectively (compared to 0.11, 0.85, and 0.95 when augmentation was enabled).
-Augmentation yielded a very small impact on performance.
-When only the last layer remained unfrozen the ML model produced a validation loss, mAP, and accuracy of 0.18, 0.87, and 0.80 respectively (compared to 0.11, 0.85, and 0.95 when all layers were trainable).
-This configuration produced the highest mAP of 0.87.
-However, a high mAP can still be achieved even when there are many false negatives in the presence of few false positives.
-The default configuration had better validation loss and accuracy results.
-Therefore, having all layers trainable yielded improved overall results.
-When using random weight initialisation, the ML model produced a validation loss, mAP, and accuracy of 0.53, 0.61, and 0.15 respectively (compared to 0.11, 0.85, and 0.95 when using weights trained on AudioSet).
-The metrics indicate the ML model produced with random weight initialisation was likely not practically useful.
-Initialising the weights, trained on AudioSet, through transfer learning was essential to achieving a performant ML model.
-The implications of the small size of AHEAD-DS can be observed in the analysis of results using random weight initialisation.
-Random weight initialisation produced a high loss for both training and validation, 0.53 and 0.53 respectively.
-This indicates underfitting.
-The learning curves show that fine tuning weights trained on AudioSet is necessary, when faced with the limited training data in AHEAD-DS.
-The tradeoff of using transfer learning is that several hyperparameters become fixed.
-Changing a hyperparameter impacts all parameters dependent on the said hyperparameter.
-A new model will need to be initialised and re-trained from scratch using AudioSet, before transfer learning can be applied.
-This makes it resource intensive to change hyperparameters such as extending the window length beyond 960ms.
-(a) Configurations vs validation loss.
-Lower values are better.
-The configuration with no label smoothing produced the lowest validation loss.
-However, this result did not lead to better mAP or accuracy.
-This occurs when predictions are wrong but still relatively close to the correct results.
-(b) Configurations vs validation mAP.
-Higher values are better.
-The configurations where only the last layer remained unfrozen produced the highest mAP.
-However, a high mAP can still be achieved even when there are many false negatives in the presence of few false positives.
-This configuration did not produce the highest accuracy and had high validation loss.
-(c) Configurations vs validation accuracy. Higher values are better. The equal best accuracy, after rounding, was produced under the default configuration.
-Figure 6: Configurations vs several validation metrics.
-The best overall result was produced under the default configuration with label smoothing enabled, augmentation enabled, all layers trainable, and weight initialisation with weights trained on AudioSet instead of random initialisation.
-6.3 Gain Sensitivity Analysis
-Tests were performed at different input gains. The validation mAP and accuracy are shown in Figure 7.
-We validated with a gain of -20 to +20 dB in 5 dB increments.
-We observed best validation mAP and accuracy results when testing with 0 dB gain.Deviations in sound levels above or below the model’s anticipated range, as determined by its training data, led to diminished performance.
-This indicates adjustment of the levels is necessary to achieve optimal performance when applying the model to new datasets.
+The regularising techniques of label smoothing and augmentation were validated.
+The model produced results which were very close to the models which did use these techniques, approximately within 0.01 mAP.
+Augmentation by applying a larger gain range of $\pm$24 dB, or use of weighted class training produced a slightly lower mAP, 0.85, but similar accuracy.
+The lowest validation loss of 0.01 was achieved when label smoothing was disabled.
+When only the last layer was trainable, the ML model produced a mean validation loss, mAP, and accuracy of 0.18, 0.91, and 0.80 respectively.
+When using random weight initialisation, the ML model produced a validation loss, mAP, and accuracy of 0.53, 0.60, and 0.15 respectively.
+Using the TAU Urban Audio Visual Scenes 2021 dataset for pre-training produced similar results to random initialisation, a mean validation loss, mAP, and accuracy of 0.53, 0.62, and 0.15 respectively.
+Normalising audio to a fixed root mean square (RMS) produced a noticeable drop in mAP and accuracy, 0.75 and 0.84 respectively.
+Similarly automatic gain control (AGC) produced a noticeable drop in mAP and accuracy, 0.77 and 0.85 respectively.
+TABLE IV: Ablation study, the mean and standard deviations for loss, mAP, and accuracy for several configurations on the validation set over 10 iterations.
+Configuration Mean Validation Loss $\pm$ SD Mean Validation mAP $\pm$ SD Mean Validation Accuracy $\pm$ SD
+Default 0.11 $\pm$ 0.00 0.88 $\pm$ 0.01 0.95 $\pm$ 0.01
+No Label Smoothing 0.01 $\pm$ 0.00 0.88 $\pm$ 0.01 0.95 $\pm$ 0.01
+No Augmentation 0.11 $\pm$ 0.00 0.87 $\pm$ 0.01 0.94 $\pm$ 0.01
+Unfreeze Only Last Layer 0.18 $\pm$ 0.01 0.91 $\pm$ 0.00 0.80 $\pm$ 0.01
+Random Weight Initialisation 0.53 $\pm$ 0.01 0.60 $\pm$ 0.00 0.15 $\pm$ 0.02
+TAU pre-training 0.53 $\pm$ 0.01 0.62 $\pm$ 0.00 0.15 $\pm$ 0.02
+Augmentation $\pm$24 dB 0.11 $\pm$ 0.00 0.85 $\pm$ 0.01 0.94 $\pm$ 0.01
+Weighted Classes 0.11 $\pm$ 0.00 0.85 $\pm$ 0.01 0.94 $\pm$ 0.01
+Normalised RMS=0.2 0.17 $\pm$ 0.01 0.75 $\pm$ 0.01 0.84 $\pm$ 0.01
+Block AGC RMS=0.2 0.17 $\pm$ 0.01 0.77 $\pm$ 0.01 0.85 $\pm$ 0.01
+III-D 2-Fold Cross Validation
+A two fold cross validation was implemented, the results are shown in Table V.
+Each fold was further divided so that ten percent of the samples within the fold were used as a hold-out dataset during training.
+This was necessary as the training loop expects the presence of a validation set.
+This created a final 45/5/50 train/validation/test split, where 45/5 belong to one fold for training and validation, and 50 belongs to the other fold for testing.
+The results show very similar performance between 2-fold cross validation and the 70/10/20 data split. Both methods achieved a loss of 0.11, mAP of 0.86, and accuracy of 0.93.
+TABLE V: The loss, mAP, and accuracy for 2-fold cross validation using AHEAD-DS.
+Fold mAP Accuracy
+Fold 1 train/Fold 2 test 0.86 0.93
+Fold 2 train/Fold 1 test 0.86 0.93
+Mean 0.86 0.93
+III-E Gain Sensitivity Analysis
+Tests were performed at different input gains.
+The validation mAP and accuracy are shown in Figure 5.
+Validation was performed with a gain of -20 to +20 dB in 5 dB increments.
+Gain was applied to both speech and background noise, so SNR did not change.Close performance was observed in validation mAP and accuracy when testing with -5, 0, and 5 dB gain.
+A higher mAP was achieved at 0 dB due to fewer false positives, but the accuracy was lower compared to -5 and 5 dB due to more false negatives.
 (a) Gain vs validation mAP. Higher values are better.
 (b) Gain vs validation accuracy. Higher values are better.
-Figure 7: The validation mAP and accuracy at different gain levels. The best results were achieved with zero gain. Higher mAP and accuracy indicate better results.
-6.4 Model Speed and Edge Deployment
-The mean processing times are shown in Figure 8.
-The tests were performed on a Google Pixel 3, a phone with modest specifications, released in 2018.
-We recorded 5, 10, 20, and 30s of audio on the phone, and measured the time it takes to process the audio by the model.
-For each length of time, we repeat the procedure five times and compute the mean time.
-We take a timestamp when the audio is buffered into memory and another timestamp when the results can be read from memory.
-The difference between the timestamps in milliseconds is recorded as the processing time.
-The processing time includes the time to compute the log mel spectrograms.
-It takes a mean time of 200.4ms to process 5s of sound.
+Figure 5: The validation mAP and accuracy at different gain levels.
+III-F Signal To Noise Ratio Analysis
+The goal of these tests were to determine the impact of the mixing and level boosts outlined in Algorithm 1. The results are shown in Table 6.
+In the original AHEAD-DS, a total of 2656 of the 6980 training files, 381 of the 1001 validation files, and 756 of the 1987 test files were processed using Algorithm 1.
+The default configuration of hyperparameters were used during training, while validation speech and environment sounds were mixed at a fixed SNRs rather than cycled through the range 10,5,0,5,10-10,-5,0,5,10 dB, both with and without the level boost.
+Best performance was achieved when the validation parameters were closest to those used in training.
+A 0 dB SNR with a level boost achieved the best performance with mAP and accuracy at 0.84 and 0.91.
+Further drops in performance were observed when the SNR was further from 0 dB.
+There was a larger drop in performance without the level boost, at 0 dB SNR the mAP and accuracy were 0.80 and 0.87.
+(a) SNR vs validation mAP. Higher values are better.
+(b) SNR vs validation accuracy. Higher values are better.
+Figure 6: The validation mAP, and accuracy for a series of SNRs.
+III-G Best Per Class Threshold
+The best per class thresholds on the validation set are shown in Table VI.
+The best threshold was calculated by finding the threshold which produced the largest value when precision is multiplied by recall.
+TABLE VI: Best per class threshold on the validation set.
+Class Threshold Precision Recall
+cocktail_party 0.41 1.00 1.00
+interfering_speakers 0.77 0.95 0.99
+in_traffic 0.56 0.97 0.97
+in_vehicle 0.58 0.98 0.99
+music 0.66 0.90 0.99
+quiet_indoors 0.35 1.00 1.00
+reverberant_environment 0.21 0.97 0.99
+wind_turbulence 0.79 0.99 0.99
+speech_in_traffic 0.71 0.95 0.90
+speech_in_vehicle 0.60 0.91 0.93
+speech_in_music 0.93 0.98 0.85
+speech_in_quiet_indoors 0.59 0.95 0.91
+speech_in_reverb_env 0.75 0.93 0.90
+speech_in_wind_turb 0.80 0.90 0.87
+III-H Context Length
+To extend the context beyond a single 960 millisecond window, results from a sequence of consecutive windows were aggregated.
+The aggregated results, as shown in Table 7, were evaluated on the validation set.
+Sequence lengths of 5, 10, 15, and 20 windows were tested.
+For aggregation, three methods were used: a simple moving average (SMA), a weighted moving average (WMA) with exponential weights, and a summation followed by a softmax function across classes.
+The exponential weights for the WMA were set as a series of powers of $e$, specifically $e^{0},e^{1},…,e^{n}$, where $n$ represents the number of windows in the sequence.
+Extending the context window beyond 960 millisecond windows improves performance.
+SMA and sum and softmax achieve the same accuracy as both methods are order-preserving functions, so the highest scoring class remains the same.
+(a) Number of windows vs validation mAP. Higher values are better.
+(b) Number of windows vs validation accuracy. Higher values are better. Simple moving average and sum and softmax produce the same accuracy.
+Figure 7: Context length: The validation mAP and accuracy using a different number of windows.
+III-I 16-Bit Model Performance and Speed on Edge Hardware
+For edge computing deployment a 16-bit model was used.
+The 16-bit model achieved a validation mAP of 0.88 and an accuracy of 0.95.
+The processing times are shown in Table VII.
+The speed tests were performed on a Google Pixel 3, a phone with modest specifications, released in 2018.
+Segments of 5, 10, 20, and 30s of audio were recorded on the phone, and the time it takes to process the audio by the model was measured.
+For each length of time, this was repeated ten times.
+Testing was not performed on streamed audio but on recorded audio.
+The demonstration application works by having a user press a button to start and then another to stop a recording.
+The application then buffers the last 5, 10, or 30 seconds of audio.
+To measure processing time, a timestamp was recorded when the audio is buffered into memory and when the results are available.
+The difference between these two timestamps, in milliseconds, is the total processing time, which includes the time required to compute the log mel spectrograms.
+It takes a mean time of 199.4ms to process 5s of sound with a SD of 10.31ms.
 There was approximately 50ms of latency to load the model, and approximately a linear increase of 30ms per 1 second of audio.
-In Figure 8, A straight line was fit to the data using linear regression.
-The line formula is $y=29.62x+48.31$.
-The Y-axis intercept is at 48.31.
-The line is almost a perfect fit to the data.
-This is a good indication that processing time increases linearly with recorded audio time.
-The latency of approximately 50ms is a one-off cost; so long as the variable representing the model remains in scope.
+A real-time factor (RTF) of between 0.038 and 0.044 was achieved over the ten iterations when processing 5s of sound.
+The RTF decreases with increasing length of processed audio.
+The latency of approximately 50ms is a one-off cost, so long as the variable representing the model remains in scope.
 The tests were performed by deploying an Android application to the test Android phone.
@@
 The user was then able to press a button and process a fixed interval of audio buffered in memory [20].
-This processing time indicates that it is feasible to create an application which is able to perform scene recognition in real time, as it takes significantly less time to process the audio compared to the audio playback time.
-There are far newer and more performant edge devices compared to our test system, so we expect lower latency and processing time with newer devices.
-Figure 8: The mean processing time for 5, 10, 20, and 30s of sound over five iterations on a Google Pixel 3.
-It takes a mean time of 200.4ms to process 5s of sound.
-There was approximately 50ms of latency to load the model, and a linear increase of 30ms per 1 second of audio.
-A straight line was fit to the data using linear regression.
-The line formula is $y=29.62x+48.31$.
-The Y-axis intercept is at 48.31.
-The line is almost a perfect fit to the data.
-This is a good indication that processing time increases linearly with recorded audio time.
-7 Discussion
-7.1 Summary of Findings
-The experiments demonstrate that YAMNet+ achieves promising performance, with a mAP of 0.83 and an accuracy of 0.93.
-The study highlights two critical findings that are essential for YAMNet+ to perform well.
-One is the necessity of transfer learning from the pre-trained YAMNet model on AudioSet to achieve satisfactory performance.
+TABLE VII: The processing times for 5, 10, 20, and 30s of sound over ten iterations on a Google Pixel 3.
+Audio length (ms) Mean Processing Time (ms) $\pm$ SD RTF Range
+5000 199.4 $\pm$ 10.31 0.037 to 0.044
+10000 344.5 $\pm$ 12.29 0.032 to 0.036
+20000 671.3 $\pm$ 22.88 0.032 to 0.035
+30000 959.8 $\pm$ 31.12 0.030 to 0.033
+III-J OpenYAMNet vs Perch vs AST
+OpenYAMNet was compared against Perch and AST.
+The results are shown in Table VIII.
+The results were produced using the test set of AHEAD-DS.
+The two models were trained and validated on the embeddings from Perch and AST using the training and validation sets of AHEAD-DS respectively.
+Since both Perch and AST require an input length of 160000 (10 seconds), the results from OpenYAMNet were aggregated using the aforementioned sum and softmax procedure, across all 20 windows in each recording.
+Furthermore, because the models were only trained on the embeddings of Perch and AST, the version of OpenYAMNet used in this test had only the last layer trainable.
+TABLE VIII: Comparison of OpenYAMNet against Perch and AST using AHEAD-DS.
+Model Parameters Test mAP Test Accuracy
+OpenYAMNet 3.7 million 0.84 0.93
+Perch 7.8 million 0.80 0.89
+AST 86.0 million 0.93 0.97
+IV Discussion
+This section discusses the results of the tests from the previous section.
+This includes the implications of the findings, a comparison to existing work, limitations of the current solution, and next steps for the future direction of the research.
+IV-A Summary of Findings
+The experiments demonstrate that OpenYAMNet achieves promising performance, with a test mAP of 0.86 and an accuracy of 0.93.
+The workflow and toolchain created for training and deploying OpenYAMNet to edge devices is useful for future research in this area.
+There were a few classes which were misidentified a notable number of times during testing.
+The class speech in music and music were often misidentified during pauses in speech and only music could be heard.
+The same issue occurs for speech in traffic and in traffic.
+The class speech in vehicle was sometimes misidentified as speech in wind turbulence because the sound of rushing air entering the vehicle can often be heard.
+The study highlights two critical findings that are essential for OpenYAMNet to perform well.
+One is the necessity of transfer learning from the pre-trained YAMNet model on AudioSet to achieve good performance.
 The other is the necessity of proper adjustment of sound levels to ensure optimal performance.
-The real-time sound-based scene recognition capabilities of YAMNet+ on edge devices were successfully demonstrated.
-On a Google Pixel 3 smartphone (a phone with modest specifications, released in 2018), the model processes audio with approximately 50ms of latency to load the model, and an approximate linear increase of 30ms per 1 second of audio.
-This indicates the feasibility of deploying such models for real-time applications on resource constrained devices.
-The workflow and toolchain created for training and deploying a scene recognition model to edge devices is useful for future research in this area.
-7.2 Comparison with Previous Work
-Compared to the original YAMNet public release, YAMNet+ is an improvement because we implemented a complete workflow, so it can be tuned with transfer learning.
-YAMNet published without training code can only be used as a black box.
-YAMNet+ is also lightweight with 3.7 million parameters compared to the >80>80 million parameters of other sound recognition models [17] [10].
-AHEAD-DS allows comparison of models using a standardised dataset with consistent labels, divided into training/validation/testing sets, and annotated with the ground truth.
-AHEAD-DS is particularly valuable given the difficulty of directly comparing existing studies, where datasets may be incomplete, not in a ready-to-use state, labels differ between datasets, and the code and datasets may not be publicly available.
-7.3 Limitations
-The testing performed in previous sections reveal several limitations of the YAMNet+ and AHEAD-DS.
-Both training and validation loss of 0.53 show the AHEAD-DS dataset is small in size, and leads to underfitting when training YAMNet+ from scratch using random weight initialisation.
-We worked around this problem by using transfer learning, where the model is initialised using weights trained on AudioSet.
-Another implication of the small dataset is the observed small improvement after applying regularisation techniques such as label smoothing and augmentation.
-A model must generalise better to perform well on a larger dataset, therefore regularisation techniques should be more impactful if the AHEAD-DS is expanded.
+Two ways of applying transfer learning were tested with OpenYAMNet.
+One method is to freeze all except the last layer, alternatively parameters in all layers are trainable.
+The Ablation Study showed that training all layers yielded better results than training only the last layer, which significantly impaired model performance.
+Another finding is that the TAU dataset is not a sufficient pre-training substitute for AudioSet.
+At approximately twelve thousand audio clips compared to the approximately two million in AudioSet, the TAU dataset is substantially smaller.
+The tradeoff of using transfer learning is that several hyperparameters become fixed.
+Changing a hyperparameter impacts all parameters dependent on the said hyperparameter.
+A new model will need to be initialised and re-trained from scratch using AudioSet, before transfer learning can be applied.
+A learning rate of 0.001 was found to perform very poorly.
+The resulting ML model was likely not practically useful.
+This was likely because its large value caused the model parameters to overshoot optimal values.
+The Ablation Study section tested several hyperparameter configurations.
+The two regularising techniques of label smoothing and augmentation did not have a major impact on the validation mAP and accuracy.
+The two techniques were kept enabled as the performance drop was minimal.
+The one noteworthy result was the validation loss of 0.01 when label smoothing was disabled.
+The result is expected as smoothing lowers the value of the true class while increasing the values of the negative classes during training.
+When a more extreme augmentation range using $\pm$24 dB was validated, there was a slight dip in performance.
+Since the gain is uniformly distributed between $\pm$24 dB, the distribution introduces occasional outliers at or near $\pm$24 dB leading to a performance drop.
+The use of weighted classes also caused a slight drop in performance.
+Focal loss already compensates for class imbalance.
+Therefore, the use of weighted classes does not seem necessary.
+AGC and normalisation caused a drop in performance.
+Both techniques discard information on the relative differences in level between environments.
+A 2-fold cross validation test was conducted to determine if the previous experiment results were an artifact of the 70/10/20 training/validation/testing split.
+The 2-fold cross validation achieved equivalent mAP and accuracy to the test set.
+This is a strong indicator that the results produced using the 70/10/20 split are not an artifact of the split.
+The 70/10/20 split was based on a rule of thumb.
+In gain sensitivity tests, deviations in sound levels outside the -5, 0, and 5 dB range led to diminished performance.
+This indicates adjustment of the levels is necessary to achieve optimal performance when applying the model to new datasets.
+In SNR analysis there was a drop in performance without the level boost from Algorithm 1.
+This was due to a large drop in metrics for speech in quiet indoors.
+Without the level boost, the addition of speech did not significantly change the signal after quantisation.
+The reason to detect speech at -10 dB SNR was to test the possibility of detecting speech, though not intelligibility, in edge cases that would challenge a human with good hearing.
+The validation results indicate it is possible to increase precision and accuracy by setting per class thresholds.
+Extending beyond a single window smooths out the results, reducing the effects of occasional errors leading to overall higher mAP and accuracy.
+When using WMA with exponential weights and a context length of 20 windows, the mAP is lower than a single window and the accuracy also drops compared to 15 windows.
+Each 10 second audio sample is divided into 20 overlapping windows, the weight of the $20th$ window is $e^{20}$, which dominates the results of the previous 19 windows.
+If the $20th$ windows is misclassified, the entire audio sample will likely be misclassified, this is detrimental to performance.
+Sum and softmax show the largest increase to mAP.
+This is because softmax further boosts the score of the class with highest confidence relative to the scores of other classes.
+The real-time sound-based scene recognition capabilities of OpenYAMNet on edge devices were successfully demonstrated on an edge device.
+With a RTF well below 1.0, this indicates the feasibility of deploying such models for real-time applications on resource constrained devices.
+There are more performant edge devices compared to our test system of a Google Pixel 3, even lower latency and processing time can be achieved with newer devices.
+The 16-Bit model deployed to the edge system achieved a 0.88 mAP, and 0.93 accuracy.
+The results indicate the 16-Bit model is practically indistinguishable from the 32-bit model.
+IV-B Comparison with Previous Work
+AHEAD-DS is based on HEAR-DS, which provides a solid foundation as the recordings of environment sounds have been open sourced with a permissive licence.
+One issue with HEAR-DS is that the dataset is not in a fixed, ready-to-use state.
+This is because some of the published data is unmixed, and some pseudo-random procedures were used during the production of the HEAR-DS dataset.
+Since the code has not been published, these procedures cannot be fully reproduced.
+The advantage of AHEAD-DS is that the dataset is in a ready-to-use state.
+This is important because it removes a variable when comparing ML models.
+The speech and environment sounds have been released in both mixed and unmixed forms, eliminating ambiguity.
+OpenYAMNet was compared against Perch and AST.
+It outperforms Perch in both performance and efficiency by achieving a higher mAP, accuracy, while using fewer parameters, 0.84, 0.93, and 3.7 million respectively compared to Perch’s 0.80, 0.89, and 7.8 million.
+Although AST surpasses OpenYAMNet with the highest mAP and accuracy at 0.93 and 0.97, it is less efficient as it uses 86.0 million parameters.
+Compared to the original YAMNet public release, OpenYAMNet is a fully open source re-implementation of YAMNet with a complete workflow, so it can be tuned with transfer learning.
+YAMNet was published without training code and can only be used as a black box.
+IV-C Limitations
+According to the authors of HEAR-DS, the 14 classes have been chosen because these classes are relevant to auditory scene recognition for hearing devices.
+These classes were chosen by audiologists and Audifon, a manufacturer of hearing aids.
+No further rationale was provided by the authors of HEAR-DS for these classes.
+Since AHEAD-DS adopts the same classes, this is a limitation inherited by AHEAD-DS.
+Compared to datasets like AudioSet, AHEAD-DS is still relatively small, which is a disadvantage.
+As shown in the Ablation Study, training from scratch using AHEAD-DS leads to suboptimal results, scoring a validation loss of 0.53.
+Transfer learning is essential, as the model performs much better with a mAP of 0.86, when fine tuning weights pre-trained on AudioSet.
 The use of transfer learning comes with its own tradeoffs.
-We are locked to classifying 960 millisecond windows, as this is a hyperparameter which is fixed by the use of weights from an existing model.
-It is possible to work around the fixed length by adding the results from multiple windows and applying a softmax mathematical operation.
-The tests of gain sensitivity analysis indicate adjustment of the levels is necessary to achieve optimal performance.
+Hyperparameters are locked in due to transfer learning.
+This makes it resource intensive to change hyperparameters such as the window length.
+Extending it beyond 960ms will require the model be trained from scratch using AudioSet.
+It is possible to work around the fixed length by aggregating the results from multiple windows and applying a softmax operation.
+The tests of gain sensitivity and SNR analysis indicate adjustment of the levels is necessary to achieve optimal performance.
 Recordings using new hardware configurations must be standardised, this only needs to be done once per device, to ensure recordings of conversations using the new device have similar levels to CHiME 6 Dev.
-The final notable limitation is the dependency on TensorFlow.
-YAMNet+ is built using the TensorFlow ML library, as it is necessary to maintain compatibility with YAMNet weights, required for transfer learning.
+A notable limitation is the dependency on TensorFlow.
+OpenYAMNet is built using the TensorFlow ML library, as it is necessary to maintain compatibility with YAMNet weights, required for transfer learning.
 TensorFlow has GPU compatibility problems, disabling the use of a GPU for quick model training and complicating the codebase.
-7.4 Next Steps
+IV-D Next Steps
 Future work can target three areas: the dataset, the model, and implementation details.
-The dataset is still relatively small compared to the likes of AudioSet and one finding is that transfer learning is essential to achieve reasonable performance.
+The dataset is still relatively small compared to the likes of AudioSet.
 Therefore, the next step is to collect more data for a new version of AHEAD-DS.
 More sound clips, more varied sound environments, and more varied recording equipment configurations ensures models trained using a future version of AHEAD-DS are more generalisable.
-We can improve the dataset by choosing scene labels in a systematic way.
-In this paper we only adopted labels from an existing study and looked at several studies which apply labels to sound clips.
-We found a lack of rigorous studies to explain why these labels are good choices.
+The dataset can be improved by choosing scene labels in a systematic way.
+In this paper class labels are adopted labels from an existing study.
+There is a lack of rigorous studies to explain why these labels are good choices.
 Examples such as speech in noise seem very generic and one talker on subway platform seem oddly specific.
-We have yet to determine the balance between labels being too specific or too generic, concerning the acoustic environments and listening contexts critical from an audiological perspective.
+The balance between labels being too specific or too generic are yet to be determined.
 This is important as good scene labels are crucial in hearing research, especially when conducting Ecological Momentary Assessment (EMA) [13].
@@
 However, if these labels are ambiguous, inconsistent, or too broad/narrow, model accuracy will suffer, negating the benefits of automation and potentially increasing the burden on participants.
-We may be able to improve the results by making better use of the limited data in AHEAD-DS.
+It may be possible to improve the results by making better use of the limited data in AHEAD-DS.
 Speech and environment sounds can be left unmixed but standardised in the training set.
-More varied training data can be generated if they are mixed on-the-fly.
-So long as the test data is not mixed on-the-fly; results still remain comparable between test runs.
-Things that would improve the model implementation would include rebuilding the model by first cleanly separating feature computation and model code.
+More varied training data can be generated if mixed on-the-fly.
+So long as the test data is not mixed on-the-fly, results still remain comparable between test runs.
+It may be possible to achieve better results by using strong annotations with AHEAD-DS. Unlike weak annotations, strong annotations precisely mark the start and end times of each sound.
+Things that would improve the model implementation include rebuilding the model by first cleanly separating feature computation and model code.
 This will enable further experimentation with other features such as time domain based features.
@@
 Newer architectures such as EfficientNet might boost precision and recall.
-We can improve the model implementation by using a modern ML library such as JAX or Pytorch.
+The model implementation can be improved by using a modern ML library such as JAX or PyTorch.
 This allows us to avoid the complex legacy code and backward compatibility problems of TensorFlow.
 It would also significantly speed up training as more modern ML libraries have better compatibility with GPU acceleration.
-We want to test the implementation on more types of edge hardware.
-This will allow us to determine the minimum resources necessary to enable real time sound recognition.
-If possible we want to deploy ML models to very power and resource constrained devices such as hearing devices and still maintain the ability to react to sounds in real time.
-8 Conclusion
+A battery and thermal impact test of the ML model on various types of edge hardware would be a valuable contribution.
+This will help determine the feasibility of using the model on more types of highly resource constrained devices such as hearing devices.
+V Conclusion
 This study addresses key challenges in scene recognition for hearing devices by introducing a refined dataset and machine learning model.
-AHEAD-DS is a publicly available dataset specifically designed for audiologically relevant environments, offering a standardised dataset for comparing models with consistent labels and in a ready-to-use state.
-Complementing this, YAMNet+ is presented as a robust sound recognition model, optimised for deployment on resource-constrained edge devices like smartphones.
-YAMNet+ only requires 3.7 million parameters compared to >80>80 million parameters of other sound recognition models.
+AHEAD-DS is a publicly available dataset repacked and refined from several other open source datasets.
+It is designed for auditory scene recognition for hearing devices and offers a standardised dataset for comparing models with consistent labels and in a ready-to-use state.
+Complementing this, OpenYAMNet is presented as a robust sound recognition model, it only requires 3.7 million parameters, and optimised for deployment on resource-constrained edge devices like smartphones.
 It serves as an open-source baseline, including model definitions, training, and testing code.
-Benchmarking YAMNet+ on AHEAD-DS yielded a mAP of 0.83 and an accuracy of 0.93.
+Benchmarking OpenYAMNet on AHEAD-DS yielded a mAP of 0.86 and an accuracy of 0.93.
 The model’s efficiency was demonstrated on a Google Pixel 3 smartphone, processing each second of audio in approximately 30ms with 50ms latency.
 The established workflow and toolchain for training and deploying such models on edge devices offer a valuable resource for future advancements in this field.
-9 Acknowledgements
-We thank our colleagues at the Macquarie University Hearing Hub and Google Research for their invaluable help, and the Google Australia Catalyst Fund and Macquarie University for their support.
-9.1 Funding
-The authors disclosed receipt of the following financial support for the research, authorship, and/or publication of this article: This work was supported by the Google DFI Catalyst fund and Macquarie University.
-9.2 Declaration of conflicting interests
-The author(s) declared no potential conflicts of interest with respect to the research, authorship, and/or publication of this article.
-9.3 ORCID iDs
-9.4 Links
-Our website and code
+Acknowledgments
+This work was supported by the Google DFI Catalyst fund and Macquarie University.